### v2

Abstract
The linear-in-means model is the standard empirical model of peer effects. Using choice data and exogenous group variation, we first develop a revealed preference style test for the linear-in-means model. This test is formulated as a linear program and can be interpreted as a no money pump condition with an additional incentive compatibility constraint. We then study the identification properties of the linear-in-means model. A key takeaway from our analysis is that there is a close relationship between the dimension of the outcome variable and the identifiability of the model. Importantly, when the outcome variable is one-dimensional, failures of identification are generic. On the other hand, when the outcome variable is multi-dimensional, we provide natural conditions under which identification is generic.
Keywords: Revealed Preference, Social Interactions, Linear-in-Means, Peer Effects
1 Introduction
People rarely make decisions in isolation and are often influenced by their neighbors and peers. [^1] Social influence has been studied in an array of contexts including test scores (Sacerdote, 2001), worker productivity (Mas and Moretti, 2009), alcohol use (Kremer and Levy, 2008), risky behavior (Card and Giuliano, 2013), and tax compliance (Fortin et al., 2007). This influence can take the form of a social norm, a common convention, or conformism. The linear-in-means model serves as a foundational framework for estimating peer effects in economics and social science research. The model supposes that an agent’s choice is a convex combination of their own ideal point and the weighted average of each other agent’s choice. The objective of this paper is to investigate the behavioral implications of the linear-in-means model leveraging revealed preference techniques à la Afriat. Moreover, we establish sufficient conditions to jointly recover each agent’s preferences and the underlying social network, including the strength of every connection.
As Manski (1993) emphatically observed, the identification of social influence parameters has long been recognized as a challenging task. The crux of the reflection problem lies in disentangling and separately identifying exogenous group effects from endogenous peer effects. In this paper, we tackle the identification of network effects through the lens of stochastic discrete choice theory—an approach that yields several novel, intuitive, and surprisingly powerful insights.
We consider a choice procedure based on the model of Manski (1993) wherein each agent chooses an action which is a convex combination of their ideal point and a weighted average of the other agents’ actions. Formally,
$p_{i}^{N}=\pi_{i}^{N}(i)v_{i}+\sum_{j\in N\setminus i}\pi_{i}^{N}(j)p_{j}^{N}$
where the vectors of $v_{i}$ and $p_{i}^{N}$ correspond the agent $i$’s ideal point and the action taken by agent $i$ in group $N$, respectively. Further, $\pi_{i}^{N}(j)\geq 0$ denotes the impact of agent $j$’s action on agent $i$ in the context of group $N$. While $p_{i}^{N}$s are observable, $v_{i}$ and $\pi_{i}^{N}$s are not observable and need to be identified from observed choices.
Consider three college friends—Ann, Ben, and Can—and their sports choices. While Can was studying abroad, Ann chose to play tennis 50% of time, volleyball 10% of the time, and walk 40% of the time, while Ben’s frequencies were 70%, 10%, and 20%, respectively. This data alone does not solve the reflection problem. An observer cannot discern whether behavioral similarities arise from peer influence or similar preferences/backgrounds that would produce identical choices even without interaction. These explanations remain observationally equivalent without further data. Our approach resolves this by leveraging natural group variation. For instance, suppose that Can returns but Ben leaves to study abroad the following semester. Now, we observe that Ann’s choices are tennis (10%), volleyball (50%), and walking (40%), while Can’s frequencies were 10%, 70%, and 20%, respectively. By comparing choice behavior across the two periods (the first semester choices of Ann and Ben without Can and the second semester choices of Ann and Can without Ben) we can uniquely isolate peer effects (point identification), provided the observed choices satisfy our model’s assumptions (formalized later). [^2] In this example, our unique identification reveals that Ann’s favorite exercise is walking, even though it is not the most frequently chosen by Ann. Indeed, her ideal point is $v_{Ann}=(0.1,0.1,0.8)$. On the other hand, tennis and volleyball are Ben’s and Can’s favorite sports, respectively. The data further show that Ann weighs her friends’ choices twice as heavily as her own ideal point (a 2:1 ratio).
[…]
Given that our identification results rely on the particular choice procedure we have adopted, it is natural to consider the falsifiability of this model. We show that our model is fully characterized by a version of classical no-money-pump conditions in the spirit of Afriat (1967). This result renders our model behaviorally testable without any restrictions across groups or about the weights assigned to each agent. Our test characterizes datasets of the previously described form which are consistent with the linear-in-means model via an easily solvable linear program. We interpret this linear program as a no money-pump condition on an outside observer who is making bets on the choices of an agent. Unlike standard no-money-pump conditions, which are typically given by two conditions, feasibility of a bet and (expected) profitability of the bet, our condition has a third part which imposes incentive compatibility of a bet. Incentive compatibility captures the idea that if the outside observer is betting on one decision maker across each group, there is no group where they would prefer to bet on a different decision maker. [^5] Testing models of peer effects is difficult as many people choose peers who are similar to them in observable characteristics. As a result of this difficulty, many studies that aim to test models of peer effects, including the linear-in-means model, do so through natural/quasi/pure experimental methods (Sacerdote, 2014; Basse et al., 2024). In addition, experimental methods are also often used when quantifying the impact of peer effects. See Agranov et al. (2021) as an example.
Thus far, we have made no assumptions about the underlying social network structure across groups. While this is a strong point of our testing and identification results, this poses a problem for the predictive power of our model. To solve this problem, we consider a refinement of the linear-in-means model, which assumes a common social network structure across groups. The strength of network connections from one person to another only varies across groups due to renormalization. Under this assumption, we provide sufficient conditions that allow us to predict choices in any possible group. To test the validity of this assumption, we develop an extension of our testing procedure for the general case. This version of the linear-in-means model is characterized by a strengthening of our previous no incentive compatible money pump condition. The updated version of incentive compatibility corresponds to the idea that if the outside observer is betting on one decision maker across each group, there is no other agent that the outside observer would rather bet on across all of the same groups.
Finally, we consider a version of the linear-in-means model more in line with the original reflection problem posed by Manski (1993). In this version of the model, each person influences each other person uniformly. This corresponds to the unweighted average choice of the group being a common social norm within the group. In this case, we provide a test in terms of a finite set of linear inequalities. In the context of linear social influence models, such as the linear-in-means model, the influence one person exerts on another is proportional to the difference in their choices. Following this logic, we call the difference between agent $i$’s choice and agent $j$’s choice the peer effect of agent $j$ on agent $i$. This version of the model is characterized by three restrictions on the peer effects between agents. Our first axiom asks that the peer effect of agent $j$ on agent $i$ is group invariant. Our second axiom is a condition about the symmetry of agent $i$’s peer effect across groups. The last axiom asks that the total peer effect on agent $i$ in group $N$ is bounded above by agent $i$’s choice in group $N$.
The rest of this paper is organized as follows. In Section 2, we formally introduce the linear-in-means model and our notation. In Section 3, we introduce and discuss our testing and identification results for the three different specifications of the linear-in-means model. In Section 4, we discuss how our results extend to alternative types of data. Finally, we conclude with a discussion of the related literature in Section 5.
2 Model and Preliminaries
Our interest is in studying the linear-in-means model of social interaction. We build on the base model in two meaningful ways. First, instead of restricting an agent’s choice to be from an interval, we allow agents to choose a distribution over a finite set of goods. Second, we consider stochastic choice data that arises when the set of agents present varies.
2.1 Preliminaries and Notation
Denote by $\mathcal{A}$ the grand set of agents. Assume that $|\mathcal{A}|\geq 2$. A typical group of agents will be denoted $N$, where $\varnothing\neq N\subseteq\mathcal{A}$. We let $\mathcal{N}\subseteq 2^{\mathcal{A}}$ be any set of groups. For any agent $i\in\mathcal{A}$, let $\mathcal{N}_{i}=\{N\in\mathcal{N}:i\in N\}$ denote the set of groups to which $i$ belongs. We sometimes abuse notation and use $N\setminus i$ to denote the group of agents formed by removing agent $i$ from group $N$. Let $X$ be some finite set of alternatives. Enumerate these alternatives from $1$ to $|X|$. We use $(0,\dots,0,1,0,\dots,0)$, where $1$ is in the $n$th dimension, to denote the $n$th alternative and use $x,y\in X$ to denote arbitrary alternatives. The data in our model consists of a probability distribution over $X$, for each $N\in\mathcal{N}$ and each $i\in N$. Formally, for $i\in N\in\mathcal{N}$, this is denoted $p_{i}^{N}\in\Delta(X)$. We use $p_{i}^{N}(x)$ to denote the choice probability of good $x$ by agent $i$ in group $N$ and $p^{N}$ to denote the matrix where each row corresponds to $p_{i}^{N}$ for a different $i\in N$. We will sometimes use $p_{-i}^{N}$ to denote the matrix which is formed by taking $p^{N}$ and removing the row corresponding to agent $i$.
2.2 The Model
In the linear-in-means model, each agent’s action is a mixture of their own ideal point and the actions of others. We use $v_{i}\in\Delta(X)$ to denote agent $i$’s ideal point. This corresponds to the action agent $i$ would take in isolation. We introduced $p_{i}^{N}$ as our data, but it also corresponds to the action taken by agent $i$ in group $N$. The amount that agent $j$’s action impacts agent $i$’s action may differ from the amount that agent $k$’s action impacts agent $i$’s action. In fact, the impact of agent $j$’s action on agent $i$ may depend on the context or the group of agents currently present. We use $\pi_{i}^{N}(j)\geq 0$ to denote the impact of agent $j$’s action on agent $i$ in the context of group $N$. Similarly, we use $\pi_{i}^{N}(i)>0$ to denote the impact of agent $i$’s ideal point on agent $i$ in group $N$. We assume that $\sum_{j\in N}\pi_{i}^{N}(j)=1$. These impact or influence weights enter into agent $i$’s action in a linear manner.
$p_{i}^{N}=\pi_{i}^{N}(i)v_{i}+\sum_{j\in N\setminus i}\pi_{i}^{N}(j)p_{j}^{N}$ (1)
Equation 1 tells us that agent $i$’s action in group $N$ is given as a convex combination of their ideal point and a weighted average of the other agents’ actions. We now discuss three key modeling assumptions which differ from the most general version of the linear-in-means model.
Our first assumption is that an agent’s outcome is restricted to the simplex. This means that the total value of agent $i$’s outcome and the total value of agent $j$’s outcome are restricted to be the same. This rules out situations where the total value of each agent’s outcome can vary, either across groups or across agents. [^6] One such example of a situation is the case of exam scores. Suppose each dimension corresponds to an agent’s score on an exam in a specific subject. If we want to allow for the sum of scores to vary across agents or groups, this simplex assumption is a meaningful restriction. In Section 4 we discuss how our main results on testing and identification extend to the case when outcomes are allowed to be anywhere in a convex subset of Euclidean space.
Our second key assumption is that $\pi_{i}^{N}\geq 0$ and $\sum_{j\in N}\pi_{i}^{N}(j)=1$. In other words, we assume that the total influence an agent faces, including their impact on themselves, is non-negative and constant across groups. [^7] This assumption on the linear-in-means model is partially due to our data assumptions. The fact that each $p_{i}^{N}$ lies in the simplex necessitates $\sum_{j\in N}\pi_{i}^{N}(j)=1$. If this were not the case, our data would lie outside the simplex. Under this assumption, the linear-in-means model is best interpreted as a model of conformism. Our agent is influenced by the weighted average of the other agents’ choices and has some incentive to match this average. In our main results, we cannot dispense with this assumption.
Our last key assumption is about the value of $\pi_{i}$ across the dimensions of the choice problem. Since we are working in a multidimensional setup, we make assumptions on our network which have no content in the one-dimensional case. In our setup, since we are considering multiple dimensions, the importance of agent $j$ to agent $i$ could in theory depend on the dimension or good in consideration. Our specification of the linear-in-means model imposes that these influence parameters, $\pi_{i}^{N}(j)$, are constant across each dimension. Our main results also rely on this assumption and it cannot be relaxed.
While this assumption is not applicable to every environment (see Belhaj and Deroïan (2014) and Zenou and Zhou (2024)), we argue that this assumption is reasonable in certain cases. As a first example, we suggest there may be certain sets of choices that share the same network. Consider the choices of professors in a department. When considering how to split their time between research, teaching, and departmental service, it is reasonable to expect that one professor’s influence on another is constant across these dimensions. As a second example, we suggest that this assumption may be more reasonable for agents in certain age ranges (or more generally, specific groups). Our prime example is the choice of children or teenagers who have yet to build distinct social networks (i.e. work colleagues, research networks, in-laws, etc.) and are mostly impacted by their school friends.
A key part of our analysis is observing data across various groups. As there are no restrictions on who is present in each of these groups, the underlying network structure can vary from group to group. This group and network variation can be modeled through different assumptions on $\pi_{i}^{N}$. Intuitively, $\pi^{N}$ is a matrix that captures weighted directed influence. By assuming that $\sum_{j\in N}\pi_{i}^{N}(j)=1$, we are assuming that $\pi^{N}$ is a stochastic matrix. When we observe group variation, each group $N$ is subject to its own $\pi^{N}$. To demonstrate the richness of our framework and motivate the analysis to follow, we discuss seven examples of linear-in-means models.
  1. 1.
    The Uniform Model (ULM): This is the simplest linear-in-means model. An agent cares about each agent (including themselves) in group $N$ equally. That is, for each $i,j\in N$,
    $\pi_{i}^{N}(j)=\frac{1}{|N|}$
    (i.e. an unweighted average). As $|N|$ grows large, each agent only cares about matching their action to the average action in the rest of the group $N$. This can be thought of as analogous to a Keynesian beauty contest (Keynes, 1937). ULM corresponds to a single underlying complete, invariant, and equal-weight network.
  2. 2.
    The Luce Model (LLM): This model captures a non-symmmetric complete network model. Each agent $i$ can be thought of as having, for each agent $j$, an invariant importance weight $w_{i}(j)$. The relative importance of agent $j$ to agent $i$ in population $N$ is simply the renormalization of each of these importance weights so that they sum to one. Formally, for each $i,j\in\mathcal{A}$, there is $w_{i}(j)$ such that for all $i\in\mathcal{A}$, $w_{i}(i)>0$, and $\sum_{j\in\mathcal{A}}w_{i}(j)=1$ such that
    $\pi_{i}^{N}(j)=\frac{w_{i}(j)}{\sum_{k\in N}w_{i}(k)}$
    More simply, LLM can be thought of as modeling the existence of a single underlying network that is group invariant. Choice in the context of group $N$ corresponds to renormalizing the fixed network weights.
  3. 3.
    The Club Model: This model aims to capture the differences in peer effects for different groups. We have in mind studying, for example, the peer effects of belonging to different clubs/cohorts/dorms. Let $P$ be a partition of $\mathcal{A}$ where $P=\{P_{1},\dots,P_{n}\}$. $P$ represents a collection of disjoint cohorts/clubs, where $P_{k}$ represents all the people in club $k$. Each agent is influenced by each club differently. Let $\alpha_{ik}$ be agent $i$’s weight for club $P_{k}$, where $\alpha_{ik}$ is a parameter capturing the popularity of club $k$ for club member $i$. Within the club, the influence is symmetric, as in ULM. [^8] One could generalize this example by assuming within-club influence as modeled in LLM. Formally, for each $i,j\in N$ such that $j\in P_{k}$,
    $\pi_{i}^{N}(j)=\frac{\alpha_{ik}}{\sum\limits_{P_{l}\cap N\neq\emptyset}\alpha_{il}}\frac{1}{|P_{k}\cap N|}$
  4. 4.
    The Friendship Model: This model allows us to capture peer effects based on a ranking of friends. The weight of each friend depends on the relative ordering of the friend in the group. For example, an agent’s best friend may be assigned the highest influence factor. Let $R_{i}$ (complete and transitive) be a ranking of the agents other than $i$, representing friendship. Agent $i$ assigns weight according to the ranking of the other agents. An agent puts a weight of $\alpha_{i}\in(0,1)$ on her ideal point, and then cares about the other agents according to their ranking. Agent $i$ assigns weight $w(j,r_{i}(j,N))$ to agent $j$ where $r_{i}(j,N)$ is the ranking of $j$ in $N$ in according to $R_{i}$. For each $i,j\in N$,
    $\pi_{i}^{N}(j)=(1-\alpha_{i})\frac{w(j,r_{i}(j,N))}{\sum_{k\in N}w(k,r_{i}(k,N))}$
  5. 5.
    The Random Participation Model: This model aims to capture the classical disconnect between analyst and agents. The analyst observes group data, such as a class roster, but is unable to observe the actual participation in that group, such as the attendance record for each agent. Each agent’s attendance is random, with the probability of agent $i$’s attending the course given by $\gamma(i)\in(0,1)$. The agents face a fixed underlying network, which is renormalized based on the realization of attendance. The weights are proportional to the amount of time spent by that agent in the same classroom. [^9] One could think of a version of the random participation model where participation is correlated across agents and is group dependent. In this case, $\phi(M,N)\in[0,1]$, with $\sum_{\emptyset\subsetneq M\subseteq N}\phi(M,N)=1$, represents the probability of $M$ being the realized group when $N$ is the observed group. This leads to the following model.$\pi_{i}^{N}(j)=\sum\limits_{i\in M\subseteq N}\phi(M,N)\frac{w_{i}(j)}{\sum\limits_{m\in M}w_{i}(m)}$.
    $\pi_{i}^{N}(j)=\sum\limits_{i\in M\subseteq N}\prod_{k\in M\setminus i}\gamma(k)\prod_{l\in N\setminus M}(1-\gamma(l))\frac{w_{i}(j)}{\sum\limits_{m\in M}w_{i}(m)}$
  6. 6.
    The Random Arrival Model: Think of a group of friends who arrive at the cafeteria at different times. This creates a different seating arrangement every time. Over the course of repeated interactions, agents face a distribution over these arrangements. The impact of one agent on another corresponds to a mixture of their impact across various fixed networks/arrangements. Let $q(A)$ be the probability of arrangement $A$.
    $\pi_{i}^{N}(j)=\sum\limits_{A}q(A)\frac{w_{A}(j)}{\sum\limits_{k\in N}w_{A}(k)}$
  7. 7.
[…]
Each model in the list above is interesting in its own right. However, due to space constraints, a comprehensive study of all of them is infeasible. Instead, we focus on the extreme cases–namely, the ULM, LLM, and GLM. Our analysis of the most general form of the GLM will provide insights into identification across these models. That said, each model benefits from additional identification power due to its specific structural assumptions. We therefore encourage future research to explore these models in greater depth.
Our goal in this paper is to study data that arises from the ULM, LLM, and GLM versions of the linear-in-means model. Our focus is on testing these models and identifying $v_{i}$ and $\pi_{N}$. With this in mind, we introduce our definition of consistency.
Definition 1.
We say that a dataset $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with GLM/LLM/ULM if there exists $v_{i}\in\Delta(X)$ for each $i\in\mathcal{A}$ and $\pi^{N}$ satisfying the conditions of GLM/LLM/ULM such that Equation 1 holds for each $N\in\mathcal{N}$ and $i\in N$.
Note that these three models are nested: ULM $\subset$ LLM $\subset$ GLM.
2.3 Data and Model Foundations
In this section, we discuss two foundations for the linear-in-means model as well as potential data-generating procedures under these foundations. In the standard setup of the linear-in-means model, when $|X|=2$, which we call the one-dimensional case as one dimension is a sufficient statistic for the other, $p_{i}^{N}$ is often interpreted as an effort level or a test score. Notably, our focus will be on situations where our data can be multidimensional. Our model allows us to capture higher granularity in choice. As we will see in Section 3, the multi-dimensional aspect of our model allows us to recover generic joint identification of preference and network parameters which generically fails in the one dimension case. [^12] As we will see later, the reflection problem of Manski (1993) is partially a result of the one dimension case being the standard case.
Our first interpretation of the linear-in-means model is as arising as the equilibrium best response of a complete information game. Consider two agents who have preferences over how they split their time between labor, leisure, and volunteering. In addition, these two agents have a desire to spend time with each other. Upon making their preferences known, the agents commit to how they spend their time. In the context of this story, an analyst has access to data on how these two agents have used their time across these three dimensions. Formally, each agent has preferences over the elements of the simplex but faces perturbations of their preferences due to the choices of their peers. [^13] In this case, an agent can be thought of as having a preference over their choice frequencies rather than repeatedly maximizing (potentially different) static utility functions. This corresponds to an agent who is deliberately stochastic (Cerreia-Vioglio et al., 2019) or faces a perturbed utility function (Fudenberg et al., 2015) subject to social influence. Specifically, the choices described in Equation 1 arise from the group $N$ playing a complete information game where each agent’s utility is given by the following. [^14] In Equation 2, note that if we choose to optimize in each dimension, ignoring across dimension constraints, we are left with $p_{i}^{N}(x)$ satisfying Equation 1. Since each dimension is subject to the same $\pi_{i}^{N}$, the proposed $p_{i}^{N}(x)$ satisfy non-negativity and $\sum_{x\in X}p_{i}^{N}(x)=1$. Notably, if we relax the assumption of $\pi_{i}^{N}(x)=\pi_{i}^{N}(y)$, allowing for non-trivial multiplexing across dimensions, our non-negativity constraints may potentially bind and our formulation would not result from optimization of a quadratic loss utility function. See Belhaj and Deroïan (2014) for a discussion of this in the two dimension case.
$u(p_{i}^{N},p_{-i}^{N})=-\pi_{i}^{N}(i)\sum_{x\in X}(p_{i}^{N}(x)-v_{i}(x))^{2}-\sum_{j\in N\setminus i}\pi_{i}^{N}(j)\sum_{x\in X}(p_{i}^{N}(x)-p_{j}^{N}(x))^{2}$ (2)
Under this interpretation, each agent’s connection weights $\pi_{i}^{N}(j)$ depend on the relative costliness of deviating from agent $j$’s choice or their own bliss point. This is in line with the observation made in Blume et al. (2015), Boucher and Fortin (2016), Kline and Tamer (2020), and Ushchev and Zenou (2020) that the one-dimensional linear-in-means model arises from agents maximizing quadratic loss functions. Under this interpretation of the model, we can think of our data as arising from each agent’s choice of time usage. We also note that Golub and Morris (2020) offers an interpretation of the linear-in-means model as arising from an incomplete information game where, in Equation 1, $v_{i}$ corresponds to agent $i$’s expectation of some underlying random variable and $p_{i}^{N}(j)$ corresponds to agent $i$’s expectation of agent $j$’s action.
Our second interpretation of the linear-in-means model is as the time average of a discrete choice procedure. Consider two agents who repeatedly go out on dinner dates with each other. Each agent has a base random utility function over the food on the menu, but this utility function is perturbed by a preference to conform with their date’s expected choice. In this story, an analyst has access to the repeated choices or the long run average of these repeated choices of each agent. Formally, in each period, each agent chooses the alternative which maximizes
$u(x)=\ln\left(\pi_{i}^{N}(i)v_{i}+\sum_{j\in N\setminus i}\pi_{i}^{N}(j)\mathbb{E}\left[\mathbf{1}\{c_{j}=x\}\right]\right)+\epsilon_{x}$ (3)
where $c_{j}$ corresponds to the choice of agent $j$ and $\mathbb{E}$ corresponds to the expected value. We make two further assumption. First, $\epsilon_{x}$ is distributed according to the Gumbel distribution, leading to logit choice frequencies (McFadden, 1974). [^15] Note that the assumption of Gumbel errors can be relaxed given that we jointly change the specification of the utility function. Second, each agent’s beliefs are stationary and correct. This means that $\mathbb{E}\left[\mathbf{1}\{c_{j}=x\}\right]$ is equal to $p_{j}^{N}(x)$. [^16] While we make no claim about convergence of such a process, this correct expectations assumption can be thought of as arising from a stationary distribution of the process where, in each period, agents choose to maximize $\ln\left(\pi_{i}^{N}(i)v_{i}+\sum_{j\in N\setminus i}\pi_{i}^{N}(j)\hat{p}_{j}^{N}(x)\right)+\epsilon_{x}$ where $\hat{p}_{j}^{N}(x)$ is the empirical frequency thus far of the choice of $x$ by agent $j$ in group N. Under these two assumptions, the long run choice frequencies are given by Equation 1.
3 Results
In this section, we study the behavioral implications of the general, the Luce, and the uniform linear-in-mean models. We begin with the general case and provide two characterizations of datasets consistent with the general case as well as partial and point identification results for $v_{i}$ and $\pi^{N}$. Our first characterization is via the non-empty intersection of a collection of convex sets. These convex sets correspond to the set of $v_{i}$ which feasibly induce the observed choice of agent $i$ the context of a given group. Our partial identification of $v_{i}$ builds on this result. Our second characterization provides an existential linear program which fails to hold if and only if the dataset is consistent with the general linear-in-means model. This linear program can be interpreted as a no money pump condition with an added incentive compatibility condition which allows for heterogeneous peer effects across groups. We then proceed to the Luce and uniform linear-in-mean models and provide refinements of these results.
3.1 The General Model
We now begin our analysis of the general linear-in-means model. Recall that in GLM, for each group of agents $N$, each agent’s choice $p_{i}^{N}$ is written as a convex combination of $v_{i}$ and $p_{j}^{N}$ across all $j\in N\setminus i$. Further, there are no restrictions across groups about the weights assigned to each agent other than each agent $i$ puts some weight on $v_{i}$. This means that testing GLM amounts to finding if there is some $v_{i}$ which can induce, for all $N\in\mathcal{N}_{i}$, $p_{i}^{N}$ as a convex combination of $v_{i}$ and $p_{j}^{N}$. This observation tells us two things. First, testing GLM can be done agent by agent. Whether or not agent $i$ has a feasible rationalizing $v_{i}$ can be tested independently of agent $j$. Second, a key part of testing GLM is finding the set of feasible $v_{i}$ for agent $i$ given $p^{N}$ for each $N\in\mathcal{N}_{i}$.
With this in mind, suppose we observe data $p^{N}$ and we see that agent $i$’s choices, $p_{i}^{N}$, lie on the interior of the convex hull of the other agents’ choices, which we write $int\Delta(p_{-i}^{N})$. In this case, no matter what $v_{i}\in\Delta(X)$ we consider, $p_{i}^{N}$ can be written as a convex combination of the other agents’ choices. Thus we can simply ask that agent $i$ puts a vanishingly small amount of weight on $v_{i}$ in their convex combination and rationalize this $v_{i}$. Now suppose we observe data such that $p_{i}^{N}$ does not lie in $int\Delta(p^{N}_{-i})$. In this case, a $v_{i}$ is feasible if and only if $p_{i}^{N}$ can be written as a convex combination of $v_{i}$ and some convex combination of $p_{j}^{N}$ for $j\in N\setminus i$. Thus the set of points on the “opposite side" of $p_{i}^{N}$ from $\Delta(p^{N}_{-i})$ correspond to the set of feasible $v_{i}$. Formally, this is given by the following equation.
$co^{-1}(\Delta(p^{N}),p_{i}^{N})=\{v\in\Delta(X)|v=\sum_{j\in N}\gamma_{j}p_{j}^{N},\gamma_{j}\leq 0\text{ }\forall j\in N\setminus i,\sum_{j\in N}\gamma_{j}=1\}$ (4)
Figure 1 visualizes the set of feasible $v_{i}$ for agent $1$ when there are three goods and three agents. These two observations tell us that the testable content of GLM amounts to checking, for each $i$, if there is some feasible $v_{i}$ which is in each $co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Figure 1: The figure considers three agents choosing when three alternatives are available. In the general linear-in-means model, $p_{1}^{N}$ is an arbitrary convex combination of $v_{1}$, $p_{2}^{N}$, and $p_{3}^{N}$. The shaded region corresponds to the set of feasible $v_{1}$ which could induce $p_{1}^{N}$ given $p_{2}^{N}$ and $p_{3}^{N}$. This corresponds to $co^{-1}(\Delta(p^{N}),p_{1}^{N})$ from Equation 4.
Our previous observations tell us that testing GLM amounts to checking if, for each $i$, a collection of convex sets has a point of mutual intersection. Keeping in line with Samet (1998) and Morris (1994), we can transform this condition into a no money-pump condition. With this in mind, we consider a setting where an outside observer is able to make bets on an agent and their choices in each group $N$.
Definition 2.
A set of vectors $\{b^{N}\}_{N\in\mathcal{N}_{i}}$ with $b^{N}\in\mathbb{R}^{X}$ for each $N\in\mathcal{N}_{i}$ is called a bet on agent i.
We restrict attention to bets which we call feasible. Effectively, feasibility is a statement about initial investment and says that, once we aggregate across each group $N$ in $\mathcal{N}_{i}$, placing a bet on agent $i$ choosing alternative $x$ should be ex-ante costly.
Definition 3.
[…]
However, for an outside observer to make a bet, it should be ex-post profitable to them. We ask that this individual rationality condition holds for each group $N\in\mathcal{N}_{i}$.
Definition 4.
A bet on agent $i$ is individually rational if $b^{N}\cdot p_{i}^{N}>0$ for each $N\in\mathcal{N}_{i}$.
We impose one last condition on these bets. Notably, we have defined these bets as bets on a specific agent $i$. To this end, we restrict to bets which are incentive compatible. Here, incentive compatibility means that the outside observer cannot gain by placing the bet on agent $j$ instead of agent $i$ at any $N\in\mathcal{N}_{i}$.
Definition 5.
A bet on agent $i$ is incentive compatible if $b^{N}\cdot(p_{i}^{N}-p_{j}^{N})\geq 0$ for each $N\in\mathcal{N}_{i}$ and each $j\in N\setminus i$.
Finally, we say that there is no incentive compatible money-pump if there is no bet that satisfies these three conditions. This effectively asks that there is no way for an outside observer to guarantee that they make an expected profit using an incentive compatible bet.
Definition 6.
We say that a dataset $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies no incentive compatible money pump if for each $i\in\mathcal{A}$ there are no strictly feasible, individually rational, and incentive compatible bets on agent $i$.
We are now ready to state our characterization of GLM.
Theorem 1.
For a dataset $\{p^{N}\}_{N\in\mathcal{N}}$, the following are equivalent.
  1. 1.
    $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with the general linear-in-means model.
  2. 2.
    For every $i\in\mathcal{A}$, the collection of sets $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ has a point of mutual intersection.
  3. 3.
    $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies no incentive compatible money pump.
We leave all proofs to the appendix. The equivalence between (1) and (2) follows from our discussion at the start of Section 3.1. If $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ corresponds to the set of feasible $v_{i}$ for agent $i$ in group $N$, then there needs to be some $v_{i}$ common to this set across all groups. The equivalence between (1) and (3) is partially a result of (2). As mentioned previously, since testing for GLM amounts to testing for a point of mutual intersection, we can transform this via linear programming duality to get our no money pump condition. However, we note that our no money pump condition does not follow immediately from (2) and an application of the result of Samet (1998) and Billot et al. (2000) as, in our proof, we work with polyhedral sets rather than compact sets. This variation allows us to recover the exact form of our no money pump condition. In Appendix D we discuss the relation and application of Samet (1998) to our Theorem 1. In this case, we get a type of no trade condition which characterizes GLM.
Before moving on, we note that condition (2) from Theorem 1 reduces to an easily checkable condition in the one-dimensional case. In the one-dimensional case, the choice frequency of one alternative is a sufficient statistic for the choice frequency of the other alternative. To simplify notation in the one dimension case, we use $p_{i}^{N}$ to denote the probability that agent $i$ chooses alternative $x$ in group $N$. Now let $\mathcal{N}_{i}^{-}\subseteq\mathcal{N}_{i}$ denote the set of groups $N$ satisfying $p_{i}^{N}\leq p_{j}^{N}$ for each $j\in N\setminus i$. Similarly, let $\mathcal{N}_{i}^{+}\subseteq\mathcal{N}_{i}$ denote the set of groups $N$ satisfying $p_{i}^{N}\geq p_{j}^{N}$ for each $j\in N\setminus i$.
Corollary 1.
In the one dimension case, a dataset $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with the general linear-in-means model if and only if, for all $i\in\mathcal{A}$, the following conditions hold.
  1. 1.
    $\min_{N\in\mathcal{N}_{i}^{-}}p_{i}^{N}\geq\max_{\mathcal{N}_{i}^{+}}p_{i}^{N}$ when both $\mathcal{N}_{i}^{-}$ and $\mathcal{N}_{i}^{+}$ are non-empty,
  2. 2.
    $\{p_{i}^{N}\}_{N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{+}}$ contains at most one value (which we denote $p_{i}^{=}$),
  3. 3.
    $\min_{N\in\mathcal{N}_{i}^{-}}p_{i}^{N}=p_{i}^{=}=\max_{\mathcal{N}_{i}^{+}}p_{i}^{N}$ when $\mathcal{N}_{i}^{-}\cap\mathcal{N}_{i}^{+}$ is non-empty.
3.1.1 Identification
We now turn our attention to the identification properties of GLM. Specifically, our focus is on the joint recovery of an agents’ bliss point as well as the network structure of each group. Going back to Manski (1993), it is known that recovering influence parameters is generally a hard problem. In this section we highlight that, conditional on observing group variation, many of these identification problems are due to the one-dimensional formulation of the original reflection problem. Our identification argument for $(v_{i},\pi_{i}^{N})$ proceeds in two steps. First, we show that, generically and with specific group variation, we are able to recover $v_{i}$ so long as $|X|\geq 3$ (i.e. in any case other than the one dimension case). Further, this identification generically fails in the one dimension case. Second, conditional on recovering $v_{i}$, we show that $\pi_{i}^{N}(j)$ is generically identified from $v_{i}$ and $p_{-i}^{N}$ so long as $|X|\geq|N|$. These two results highlight that multiple dimensions are useful in recovering each agent’s bliss point, $v_{i}$, as well as their peer influence parameters, $\pi_{i}^{N}$. We begin by discussing the partial identification of $v_{i}$ that arises without sufficient group variation.
Definition 7.
For parameter $v_{i}$, we say that $A$ is the sharp identified set if $v_{i}\in A$ if and only if there exist $\pi_{i}^{N}$ for each $N\in\mathcal{N}_{i}$ such that $p_{i}^{N}=\pi_{i}^{N}(i)v_{i}+\sum_{j\in N\setminus i}\pi_{i}^{N}(j)p_{j}^{N}$.
Our first goal is to characterize the sharp identified set of $v_{i}$. Recall our discussion of $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ prior to Theorem 1; it corresponds to the set of feasible $v_{i}$ given observed choices $p^{N}$. Theorem 1 tells us that our dataset is consistent with GLM if and only if, for each $i\in\mathcal{A}$, $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ have a point of mutual intersection. This amounts to testing if there is some ideal point $v_{i}$ which is feasible in each group $N\in\mathcal{N}_{i}$. It follows from similar logic that any point in the mutual intersection of $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ is a feasible value for $v_{i}$.
Proposition 1.
In GLM, the sharp identified set for $v_{i}$ is given by $\bigcap_{N\in\mathcal{N}_{i}}co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Proposition 1 formally states the observation from the previous paragraph. As mentioned earlier, one of our main goals is to give conditions under which $v_{i}$ and $\pi_{i}^{N}$ are jointly point identified. Here, point identified means that the sharp identified set is a singleton. As a first step in our two step procedure, our next result gives a sufficient condition for point identification of $v_{i}$. Let $\mathcal{N}_{i}^{ext}\subseteq\mathcal{N}_{i}$ denote the set of groups $N$ with $p_{i}^{N}\not\in\Delta(p_{-i}^{N})$.
Corollary 2.
Let $N_{j}=\{i,j\}$ and $N_{k}=\{i,k\}$. Suppose that $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with GLM, $N_{j},N_{k}\in\mathcal{N}_{i}^{ext}$, and that the vectors $(p_{i}^{N_{j}}-p_{j}^{N_{j}})$ and $(p_{i}^{N_{k}}-p_{k}^{N_{k}})$ are linearly independent. Then $v_{i}$ is point identified.
Figure 2: The figure considers choice by agents $1$, $2$, and $3$. We observe choice in two groups; $N_{2}=\{1,2\}$ and $N_{3}=\{1,3\}$. Consider the two rays which trace out $co^{-1}(\Delta(p^{N}),p_{1}^{N})$ for each group. These rays intersect, so, by Theorem 1, this dataset is consistent with GLM. The linear independence of $(p_{1}^{N_{2}}-p_{2}^{N_{2}})$ and $(p_{1}^{N_{3}}-p_{3}^{N_{3}})$ gives us that these rays intersect at a single point. This single point corresponds to the uniquely feasible $v_{1}$.
Figure 2 offers a visualization of Corollary 2. While we do not formally show it here, Corollary 2 naturally extends. Consider $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ for a group of agents $N$. Note that, in the case of $N\in\mathcal{N}_{i}^{ext}$, if we drop the non-negativity restriction on $v$ in our definition of $co^{-1}(\Delta(p^{N}),p_{i}^{N})$, then $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ defines a polyhedral cone. The extremal rays of this cone take the form $p_{i}^{N}-p_{j}^{N}$ plus some location translation. In out setting, linear independence of the extremal rays of two of these $n$ dimensional cones will make the intersection of these two sets $(n-1)$ dimensions. Thus, when we observe $n$ groups of $n$ agents, with each group being contained in $\mathcal{N}_{i}^{ext}$, and the set of extremal rays of $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ across all $n$ groups of $n$ agents being linearly independent, $\bigcap_{N\in\mathcal{N}_{i}^{ext}}\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}$ is a $(n-n)=0$ dimensional set. By Theorem 1 this set is non-empty and thus we get point identification. We now return to our example with Ann, Ben, and Can to highlight Corollary 2.
Example 1.
Recall the choices of Ann, Ben, and Can. While Can is studying abroad, Ann chooses $(0.5,0.1,0.4)$, (tennis, volleyball, walking), and Ben chooses $(0.7,0.1,0.2)$. While Ben is studying abroad, Ann chooses $(0.1,0.5,0.4)$ and Can chooses $(0.1,0.7,0.2)$. We now calculate the difference between Ann’s choices and Ben’ and Can’s choices in their respective groups. When both Ann and Ben are choosing, the difference in their choices is given by $(0.2,0,-0.2)$. Similarly, when both Ann and Can are choosing, the difference between their choices is given by $(0,0.2,-0.2)$. As these two vectors are linearly independent, Corollary 2 tells us that we can point identify Ann’s ideal point. Ann’s ideal point is given by the point of intersection between the two lines $(0.5,0.1,0.4)+t(0.2,0,-0.2)$ and $(0.1,0.5,0.4)+r(0,0.2,-0.2)$. The intersection of these two lines is guaranteed by Theorem 1. Calculating the point of intersection gives us $(0.1,0.1,0.8)$, thus identifying Ann’s ideal point. This also tells us that, even though walking is never Ann’s most chosen activity, it is her most preferred activity in isolation.
Before moving on, we point out that point identification of $v_{i}$ generically fails in the one-dimensional case. Recall the three conditions of Corollary 1. The last two conditions arise as a result of non-generic situations. Specifically, if we observe data where $p_{i}^{N}\neq p_{i}^{M}$ for all $N,M\in\mathcal{N}_{i}$, then it is sufficient to just check the first condition. However, in this case, $v_{i}$ is necessarily not point identified. In the one-dimensional case, the sharp identified set corresponds to the interval $[\max_{\mathcal{N}_{i}^{+}}p_{i}^{N},\min_{N\in\mathcal{N}_{i}^{-}}p_{i}^{N}]$ and this set is singleton only in a non-generic case. We visualize this failure of point identification in Figure 3 and show it in Example 2.
Figure 3: The figure considers choice by agents $1$, $2$, and $3$. We observe choice in two groups; $N_{2}=\{1,2\}$ and $N_{3}=\{1,3\}$. Unlike in Figure 2, $v_{1}$ is not point identified here. This is due to the fact that, in the one dimension case, $(p_{1}^{N_{2}}-p_{2}^{N_{2}})$ and $(p_{1}^{N_{3}}-p_{3}^{N_{3}})$ will always be linearly dependent; they will be some multiple of $(1,-1)$. Geometrically, the rays corresponding to $co^{-1}(\Delta(p^{N}),p_{1}^{N})$ for $N_{2}$ and $N_{3}$ can intersect at a single point if and only if $p_{1}^{N_{2}}=p_{1}^{N_{3}}$.
Example 2.
Consider a counterfactual world with Ann, Ben, and Can where their college no longer has tennis courts. In this world, Ann, Ben, and Can have the choice of splitting their exercise time between volleyball and walking. While Can is studying abroad, Ann’s choices are given by $(0.8,0.2)$, (volleyball, walking), and Ben’s choices are given by $(0.9,0.1)$. While Ben is studying abroad, Ann’s choices are given by $(0.7,0.3)$ and Can’s choices are given by $(0.85,0.15)$. When both Ann and Ben are choosing, the difference in their choices is given by $(-0.1,0.1)$. On the other hand, when both Ann and Can are choosing, the difference in their choices is given by $(-0.15,0.15)$. Notice that these two vectors are multiples of each other and are thus linearly dependent. These choices, as would any other choices in this two-alternative world, fail the conditions of our Corollary 2. Ann’s choices when Ben is abroad are given by $(0.7,0.3)$ and, for that to lie in the convex hull of $(\alpha,1-\alpha)$ and $(0.85,0.15)$, $\alpha$ must be weakly less than $0.7$. Similarly, her choices when Can is abroad are given by $(0.8,0.2)$ and, for this to be between $(0.9,0.1)$ and $(\alpha,1-\alpha)$, $\alpha$ must be weakly less than $0.8$. Combining our two constraints gives us that any vector $(\alpha,1-\alpha)$ which satisfies $0\leq\alpha\leq 0.7$ is a potential ideal point for Ann, hence we cannot point identify Ann’s ideal point.
We now turn our attention to the recovery of $\pi_{i}^{N}$. This recovery utilizes knowledge of $v_{i}$ and thus is dependent on our previous arguments in Corollary 2.
Proposition 2.
Suppose that $v_{i}$ is point identified. Then $\pi_{i}^{N}$ is point identified if the set of vectors including $v_{i}$ and $\{p_{j}^{N}\}_{j\in N\setminus i}$ is affinely independent. [^18] A set of points $\{p_{j}\}$ is affinely independent if for each $i$, there is no collection $\{\lambda_{j}\}_{j\neq i}$ for which $\sum_{j\neq i}\lambda_{j}=1$ where $p_{i}=\sum_{j\neq i}\lambda_{j}p_{j}$. When each point $p_{j}$ satisfies $p_{j}\geq 0$ and $\sum_{x\in X}p_{j}(x)=1$, affine independence and linear independence are equivalent.
The identification of $\pi_{i}^{N}$ follows from basic properties of simplices. Specifically, any point in the convex hull of a simplex can be written as a convex combination, with unique weights, of the extreme points of the simplex. The affine independence condition ensures that $v_{i}$ and $\{p_{j}^{N}\}_{j\in N\setminus i}$ form the extreme points of a simplex. We point out that affine independence of $v_{i}$ and $\{p_{j}^{N}\}_{j\in N\setminus i}$ puts joint restrictions on the number of agents and alternatives. Notably, when $|N|\leq|X|$, the set of points given by $v_{i}$ and $\{p_{j}^{N}\}_{j\in N\setminus i}$ is generically affinely independent. On the other hand, if $|N|>|X|$, the set of points $v_{i}$ and $\{p_{j}^{N}\}_{j\in N\setminus i}$ is never affinely independent. In terms of dimension counting, in group $N$, there are $|N|$ agents each with $|N|-1$ unknowns giving a total of $|N|(|N|-1)$ unknowns in $\pi^{N}$. Each agent has $|X|-1$ observables per group, and so the total number of observables in group $N$ is given by $|N|(|X|-1)$. This tells us that we have more observables than unknowns if $|X|\geq|N|$. In order to highlight this result, we return to Ann, Ben, and Can.
Example 3.
In Example 1, we were able to identify Ann’s ideal point as $(0.1,0.1,0.8)$. During the semester when Can is abroad, Ben’s choices correspond to $(0.7,0.1,0.2)$. Since Ann’s ideal point and Ben’s choices are distinct points, they are trivially affinely independent. By Proposition 2, this means that we are able to recover both the impact of Ben’s choices on Ann’s choices and the impact of Ann’s own ideal point on her choices. This problem amounts to solving for $\pi_{Ann}$ in the equation $(0.5,0.1,0.4)=\pi_{Ann}(0.1,0.1,0.8)+(1-\pi_{Ann})(0.7,0.1,0.2)$. Doing so gives us that $\pi_{Ann}$ is equal to $\frac{1}{3}$. Thus the relative importance of Ann’s ideal point is $\frac{1}{3}$ and the relative importance of Ben’s choices is $\frac{2}{3}$. Due to the symmetry of the problem, these weights are the same when considering the choices of Ann and Can.
We conclude our discussion of identification by putting our results in the context of the reflection problem of Manski (1993). Our focus is on a related but slightly different version of the linear-in-means model proposed in Manski (1993). First, our version of the model is primarily concerned with conformism ($\pi_{i}^{N}\geq 0$ and $\sum_{j\in N}\pi_{i}^{N}(j)=1$). Second, we allow for weighted averages (from unknown networks) while the version of Manski (1993) assumes an unweighted average across peers. We make weaker assumptions in one direction and stronger assumptions in other directions. Manski (1993) shows that identification fails in his setup. This continues to be true even if our conformism assumptions are applied to the setup of Manski (1993), primarily due to the fact that the reflection problem considers one-dimensional outcome variables. Our results in this section show how utilizing multidimensional outcome variables, along with group variation, allows the analyst to jointly recover agents’ bliss points and the underlying network structure. We take this to be prescriptive. While outcome variables can sometimes be sufficiently summarized in a one-dimensional statistic, such as labor time, by gathering more granular data and considering a higher dimensional outcome variable, we can actually recover and quantify heterogeneous peer effects (subject to the conditions of our results).
3.2 Luce Linear-in-Means
As we have seen, there are conditions in GLM under which we are able to recover both an agent’s ideal point $v_{i}$ and their social interaction parameters $\pi_{i}^{N}$ for groups that we observe. However, these identified parameters have little predictive power for groups we do not observe in GLM. Since GLM allows arbitrary variation in $\pi_{i}^{N}$ across groups, the most we can predict for agent $i$’s choice in an unobserved group $N$ is that it lies within the convex hull of $\{v_{j}\}_{j\in N}$. Our goal in this section is to consider the Luce linear-in-means model which allows us to connect our social interaction parameters across groups. LLM allows analysts to predict choices in unobserved groups subject to the conditions for identification from Section 3.1.1 holding.
Recall that LLM supposes that each agent $i$ has a weighting function, $w_{i}(j)$, which corresponds to the absolute importance of agent $j$ to agent $i$. In a group $N$, the relative importance of agent $j$ to agent $i$ is given by the renormalization of this weighting function, $\pi_{i}^{N}(j)=\frac{w_{i}(j)}{\sum_{k\in N}w_{i}(k)}$. The interpretation here is that there is a single underlying network fixed across each group. We first discuss the identification and predictive properties of LLM. We focus on the case when $v_{i}$ is known. For a specific group $N$, if the conditions of Proposition 2 hold, we can pin down $\pi_{i}^{N}$. This allows us to pin down the relative weights of each agent $j\in N\setminus i$ as $\frac{w_{i}(j)}{w_{i}(k)}=\frac{\pi_{i}^{N}(j)}{\pi_{i}^{N}(k)}$. This is the exact condition for identification that is used in the Luce model of Luce (1959). This means that if $N=\mathcal{A}$, we can predict choice (for agent $i$) on every possible group of agents. However, we may not always observe choice on $N=\mathcal{A}$ or, if we do, the conditions of Proposition 2 may not be satisfied at $N=\mathcal{A}$. To regain full predictive power in LLM, we need the conditions of Proposition 2 to hold for a collection of groups $\{N_{l}\}_{l=1}^{L}$ such that every two agents $k$ and $j$ can be compared through a string of groups.
Definition 8.
Suppose $v_{i}$ is point identified. Agents $j$ and $k$ are comparable by agent i if there exist two observed groups $N_{1},N_{2}\in\mathcal{N}$ such that the following hold.
  • $i\in N_{1}\cap N_{2}$,
  • $j\in N_{1}$ and $k\in N_{2}$,
  • For each $l\in\{1,2\}$, $v_{i}$ and $\{p_{m}^{N_{l}}\}_{m\in N_{l}\setminus i}$ are affinely independent.
Proposition 3.
Suppose that $v_{i}$ is point identified. Then $w_{i}(j)$ is point identified for all $j\in\mathcal{A}$ if every pair of agents $j\neq k$ are comparable by agent $i$.
Proposition 3 is an immediate result of our Proposition 2, Corollary 4 of Alós-Ferrer and Mihm (2024), and the fact that $i$ is in both $N_{1}$ and $N_{2}$. Further, Proposition 3 gives conditions under which we can predict the choice of agent $i$ in any group $N$, conditional on knowing $v_{j}$ for each $j\in N$. We show how to operationalize this predictive power by revisiting Ann, Ben, and Can one last time.
Example 4.
Recall the choices of Ann, Ben, and Can described in Example 1. Our goal now is to predict how Ann, Ben, and Can will choose when all three of them are choosing together. As we showed in Example 3, the weight Ann puts on Ben or Can is twice the weight she puts on her own ideal point. This tells us that Ann’s $w$ vector is given by $(0.2,0.4,0.4)$. Since we have only observed Ben and Can choosing in a single group each, we are unable to pin down their ideal points or network weights. Now suppose we observe choice from a third semester when Ann chooses to study abroad. In this semester, both Ben and Can choose $(0.5,0.5,0)$. Since neither Ben nor Can decide to walk as a form of exercise this semester, it reveals that both of their ideal points put no weight on walking. Using this observation and similar techniques to those in Example 1, we can conclude that Ben’s and Can’s ideal points are $(0.9,0.1,0)$ and $(0.1,0.9,0)$ respectively. By applying Proposition 2 and 3, we recover that Ben and Can put the same amount of weight on their own ideal points and the choices of each other agent. This means that both of their $w$ vectors are $(\frac{1}{3},\frac{1}{3},\frac{1}{3})$. Since we have recovered the ideal points and the $w$ vectors of Ann, Ben, and Can, we can now predict their choices in the semester where all three of them are choosing together. In this semester, Ann’s choices are $(\frac{11}{30},\frac{11}{30},\frac{4}{15})$, Ben’s choices are $(\frac{8}{15},\frac{1}{3},\frac{2}{15})$, and Can’s choices are $(\frac{1}{3},\frac{8}{15},\frac{2}{15})$.
Figure 4 compares GLM and LLM in terms of their prediction power on the Marschak-Machina triangle in a domain of three alternatives, $X=\{x,y,z\}$ and three agents $\{1,2,3\}$. This is the simplest environment in which we can illustrate the differences between these models. In the figure, stochastic choices for three binary groups ($\{1,2\},\{1,3\}$ and $\{2,3\}$) are fixed and denoted by different colored dots on the edges of the dotted triangle. The blue dots indicate agent $3$’s choices, $p_{3}^{\{1,3\}}$ and $p_{3}^{\{2,3\}}$. We then illustrate the predictions for $p_{3}^{\{1,2,3\}}$ for GLM and LLM. The blue-shaded triangle identifies all possibilities for $p_{3}^{\{1,2,3\}}$ for GLM given these binary group choices. This figure illustrates both the predictive and explanatory power of GLM: If $p_{3}^{\{1,2,3\}}$ lies outside this triangle, the data cannot be explained by GLM.
LLM restricts choices for $p_{3}^{\{1,2,3\}}$ even further. Indeed, LLM predicts that there is only a single possibility for $p_{3}^{\{1,2,3\}}$ indicated by a blue dot inside GLM’s prediction. If $p_{3}^{\{1,2,3\}}$ is not equal to this point, then the data cannot be explained by the Luce linear-in-means model. Note that the prediction of LLM belongs to the shaded triangle indicating that LLM is a special case of GLM.
Figure 4: This figure illustrates predictions of GLM and LLM for three alternatives, $X=\{x,y,z\}$ and three agents: blue, green, and orange. The blue-shaded triangle identifies all possibilities for the choice probability of the blue agent consistent with GLM in the entire group, given three binary group choices for each agent depicted on three edges of the shaded region. Note that LLM predicts uniquely the choice probability of the blue agent.
While we focus on LLM to recover predictive power in the linear-in-means model, one could instead consider any known mapping from $\pi_{i}^{N}$ to $\pi_{i}^{M}$ for two different groups $N$ and $M$ to recover predictive power. We focus on LLM for two reasons. First, LLM is an intuitive criterion that captures the case of there being a fixed underlying network. Second, as we now discuss, we can test LLM using a natural extension of our no incentive compatible money pump condition. Recall our story about an outside observer making bets on an agent and their choices. In the case of GLM, we asked that these bets were strictly feasible, individually rational for each group containing $i$, and incentive compatible for each group containing $i$. When moving from GLM to LLM, we are moving to a model where the social influence parameters $\pi_{i}^{N}$ and $\pi_{i}^{M}$ are actually connected across groups. Our condition for testing LLM extends the no incentive compatible money pump condition taking this across group connection into account. Specifically, we weaken incentive compatibility so that the outside observer cannot gain by placing a bet on $j$ instead of $i$ across every $N\in N_{i}\cap N_{j}$.
Definition 9.
A bet on agent $i$ is weakly incentive compatible if, for each $j\neq i$, $\sum_{N\in\mathcal{N}_{i}\cap\mathcal{N}_{j}}b^{N}\cdot(p_{i}^{N}-p_{j}^{N})\geq 0$.
Definition 10.
We say that a dataset $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies no weakly incentive compatible money pump if for each $i\in\mathcal{A}$ there are no strictly feasible, individually rational, and weakly incentive compatible bets on agent $i$.
Theorem 2.
For a dataset $\{p^{N}\}_{N\in\mathcal{N}}$, the following are equivalent.
  1. 1.
    $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with the Luce linear-in-means model.
  2. 2.
    $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies no weakly incentive compatible money pump.
Observe that if our dataset satisfies no weakly incentive compatible money pump, then it satisfies no incentive compatible money pump. It then follows from Theorem 1 that the collection of sets $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ have a point of mutual intersection. In the case of GLM, this is the set of feasible $v_{i}$. The strengthening of incentive compatibility to weak incentive compatibility is exactly what guarantees us that the social influence parameters $\pi_{i}^{N}$ follow a Luce rule across groups.
3.3 Uniform Model
In the previous sections, we considered the linear-in-means model allowing for heterogeneous social interaction terms. The interpretation of the heterogeneity is that the importance of agent $j$ and agent $k$ may differ to agent $i$. In this section, we consider the hypothesis that agent $i$ belonging to a group impacts their behavior but no one agent in that group is any more important than the other. We model this hypothesis through the uniform linear-in-means model. This is a special case of LLM when $w_{i}(j)=w_{i}(k)$ for each $j,k\in\mathcal{A}$. In this section, we maintain the assumption that each observed group has the same size, for all $N,M\in\mathcal{N}$, $|N|=|M|$. This is done so we can focus on the empirical content of variation of group make-up rather than group size. We relax this assumption in Appendix E where we give all the axioms and results from this section allowing for group size variation.
In the context of a linear influence model, such as ours, $p_{i}^{N}-p_{j}^{N}$ corresponds to a (rescaling) of the influence agent $j$ has on agent $i$’s choice. Consequently, we call $p_{i}^{N}-p_{j}^{N}$ the peer effect of agent $j$ on agent $i$ in group $N$. All of our axioms for ULM are stated in terms of peer effects. Before stating our axioms, we need one definition.
Definition 11.
A cycle is a sequence of tuples $\{(i_{k},j_{k},N_{k})\}_{k=1}^{K}$ such that $i_{k},j_{k}\in N_{k}$, $j_{k}=i_{k+1}$, and $j_{K}=i_{1}$.
A cycle captures cycles of influence. Agent $i_{1}$ influences agent $j_{1}$ in group $N_{1}$ and then agent $j_{1}=i_{2}$ influences agent $j_{2}$ in group $N_{2}$. This is repeated until we return to agent $i_{1}$. Our first axiom puts restrictions on the sum of peer effects across cycles.
Axiom 1 (Cyclically constant).
A dataset $\{p_{N}\}_{N\in\mathcal{N}}$ is cyclically constant if for every cycle $\{(i_{k},j_{k},N_{k})\}_{k=1}^{K}$ we have that
$\sum_{k=1}^{K}(p_{i_{k}}^{N_{k}}-p_{j_{k}}^{N_{k}})=0.$ (5)
To best understand cyclic constancy, we first introduce a second axiom implied by cyclic constancy.
Axiom 2 (Constant Peer Effects).
A dataset $\{p_{N}\}_{N\in\mathcal{N}}$ satisfies constant peer effects if for each $N$ and $M$ with $i,j\in N\cap M$ we have that $[p_{i}^{N}-p_{j}^{N}]=[p_{i}^{M}-p_{j}^{M}]$.
Observe that constant peer effects is implied by cyclic constancy when we consider cycles of length two. Constant peer effects tells us that the peer effect of agent $j$ on agent $i$ is group invariant. Returning to cyclic constancy, it says that the peer effect of agent $i$ on agent $j$ plus the peer effect of agent $j$ on agent $k$ should be equal to the peer effect of agent $i$ on agent $k$ (and so on for longer cycles). Further, there is no way to break this equality by going to different groups during a cycle. In this manner, cyclic constancy tells us that agent $i$ to agent $j$ peer effects are group invariant and that a long chain of peer effects corresponding to the indirect peer effect of agent $i$ on agent $j$ equals the direct peer effect of agent $i$ on agent $j$. Our next axiom puts restrictions on the peer effect of agent $i$ across different groups.
Axiom 3 (Symmetric Peer Effects).
For $i\in N\cap M$,
$p_{i}^{N}-\sum_{j\in N\setminus M}[p_{j}^{N}-p_{i}^{N}]=p_{i}^{M}-\sum_{k\in M\setminus N}[p_{j}^{M}-p_{i}^{M}].$ (6)
Symmetric peer effects tell us that, once we take agent $i$’s actual choice into account, the total peer effect of agent $i$ on agents in group $N$ which are not in $M$ is equal to the total peer effect of agent $i$ on agents in group $M$ which are not in group $N$. The combination of cyclic constancy and symmetric peer effects tells us that the total peer effect of agent $i$ is the same in groups $N$ and $M$, once we take into account their actual choice in each group. Our last axiom restricts the total amount of peer effect that agent $i$ has in a group.
Axiom 4 (Bounded Total Peer Effects).
For $i\in N$,
$p_{i}^{N}\geq\sum_{j\in N\setminus i}[p_{j}^{N}-p_{i}^{N}].$ (7)
Bounded total peer effects tell us that the total amount of peer effect agent $i$ has in group $N$ can be no more than their actual choice $p_{i}^{N}$. With this in mind, we are now ready to give our characterization of ULM. Recall that this result assumes that $|N|=|M|$ and that this assumption is relaxed in the appendix.
Theorem 3.
For a dataset $\{p^{N}\}_{N\in\mathcal{N}}$, the following are equivalent.
  1. 1.
    $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with the uniform linear-in-means model.
  2. 2.
    $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies cyclic constancy, symmetric peer effects, and bounded total peer effects.
We first note that, in Theorem 3, we can replace cyclic constancy with constant peer effects and the equivalence holds. Our focus on cyclic constancy is due to the following discussion. Consider the following the equation.
$p_{i}^{N}=\frac{1}{|N|}\hat{v}_{i}+\sum_{j\in N\setminus i}\frac{1}{|N|}p_{j}^{N}+O^{N}$ (8)
In Equation 8, $\hat{v}_{i}$ is not restricted to have non-negative elements (but still satisfies $\sum_{x\in X}v_{i}(x)=1$) and we ask $\sum_{x\in X}O^{N}(x)=0$. Choice induced by Equation 8 differs from ULM in two important ways. First, an agent’s ideal point $\hat{v}_{i}$ no longer needs to lie within the simplex. Second, every agent $i$ in group $N$ is subject to some group specific shock to tastes given by $O^{N}$. Under one additional assumption on $\mathcal{N}$, we show in Appendix E that choice according to Equation 8 is characterized by cyclic constancy. The addition of symmetric peer effects is exactly what rules out group specific shocks, reducing $O^{N}$ to be the zero vector across all $N\in\mathcal{N}$. Finally, bounded total peer effects is what induces each agent’s ideal point to be within the simplex.
We conclude our analysis with the observation that ULM is trivially point identified. Since each agent’s social interaction parameters are pinned down to be $\frac{1}{|N|}$, the only parameter left to identify is each agent’s ideal point. This can be easily recovered from the following equation.
$v_{i}=|N|p_{i}^{N}-\sum_{j\in N\setminus i}p_{j}^{N}$ (9)
Equation 9 also offers an alternative characterization of ULM. A dataset is consistent with ULM if and only if the value of the right hand side of Equation 9 is group invariant and lies in the simplex.
4 Extensions
In this section, we discuss how our results from Section 3 extend under alternative assumptions on the data. Our first and main extension is to consider relaxing the assumption that our data satisfies $\sum_{x\in X}p_{i}^{N}(x)=1$ (as well as $p_{i}^{N}\geq 0$). In this case, all of our results extend (mostly) unchanged. We then consider three additional forms of variation and show how they can provide even stronger results than what we have discussed in Section 3.
4.1 Choices in Euclidean Space
Thus far, we have taken the perspective of stochastic choice. Agents in our model choose a probability distribution. However, nothing precludes us from investigating more sophisticated choice environments. The classic one-dimensional framework has as its choice space $[0,1]$. The unit interval can be interpreted in many ways—any cardinal unidimensional choice can be modeled this way. We have interpreted the unit interval as a probability distribution over two objects, leading to the simplex as the natural generalization. Similarly, if we were to imagine an individual deciding how to spend their day, $[0,1]$ might measure the proportion of hours devoted to sleep vs. waking, whereas a model with three alternatives might further refine waking hours into work and leisure.
However, suppose that $[0,1]$ represents the score on a test. We might also be interested in scores on multiple exams: a member of $[0,1]^{2}$ might reflect the scores on an English language exam and on a Turkish language exam. In fact, Equation (1) is meaningful whenever there is a convex subset of a vector space $Y$, and each $p_{i}^{N}$ lies in $Y$. The particular case of $Y=\Delta(X)$ is only one example. All three of our particular models (general, Luce, and uniform) also have ready interpretation in terms of more general vector spaces.
Let us illustrate this point via example, whereby our choice space is $Y=[0,1]^{T}$. We can think of $T$ as indexing a collection of tests. Any $p_{i}^{N}\in Y$ is a vector of test scores. Our equation $p_{i}^{N}=\pi_{i}^{N}(i)v_{i}+\sum_{J\in j\in N\setminus i}\pi_{i}^{N}(j)p_{j}^{N}$ asserts that an individual’s test scores are an average of some “baseline scores” ($v_{i}$) and the scores of other individuals present.
In this more general setup, the definition of inverse cone remains unchanged and still captures the set of feasible $v_{i}$ given the data.
$co^{-1}(Y,p_{i}^{N})=\{v\in Y|v=\sum_{j\in N}\gamma_{j}p_{j}^{N},\gamma_{j}\leq 0\text{ }\forall j\in N\setminus i,\sum_{j\in N}\gamma_{j}=1\}$ (10)
Theorem 1 parts 1 and 2 remain equivalent. More generally, the duality conditions (such as incentive compatible money pump) will change from domain to domain, but can be characterized for polyhedral domains. The downside is that, outside of the simplex domain, these dual conditions become more difficult to interpret. In regards to identification, since $co^{-1}(Y,p_{i}^{N})$ still captures the set of feasible $v_{i}$, all of the results from Section 3.1.1 go through unchanged.
One remaining caveat is that, in more general vector spaces, we may allow for the possibility that individual weights need not sum to one. As mentioned in Section 2, this causes problems for many of our results. For example, if $Y=\mathbb{R}^{M}_{+}$, then it is mathematically meaningful to allow any weights $\pi_{i}^{N}$ for which $\pi_{i}^{N}(j)\geq 0$ for all $j$ and $\pi_{i}^{N}(i)>0$. Which sets of weights might be mathematically meaningful depends, in general, on the domain under consideration. The economic interpretation will likely dictate which set of weights to choose.
4.2 Other Forms of Variation
Throughout this paper, we have focused on group variation being the main source of variation in the data. We now consider three alternative form of variation; network variation, characteristic variation, and product-attribute variation. We discuss each of these forms of variation in a series of three remarks.
Remark 1 (Network Variation).
Instead of (or in addition to) group variation, we could consider variation of the underlying network. That is to say, fix a group $N$ and consider two realizations of some observable variable, such as time period, or regime, $R$ and $R^{\prime}$. Suppose that an agent’s social influence parameters are given by $\pi_{i}$ when $R$ is observed and $\pi^{\prime}_{i}$ when $R^{\prime}$ is observed. Neither $\pi_{i}$ nor $\pi^{\prime}_{i}$ are observable to the analyst. If no restrictions are placed on $\pi_{i}$ or $\pi^{\prime}_{i}$, testing for the linear-in-means model with network variation can be done using our Theorem 1. Further, all of the identification results of Section 3.1.1 also extend to this arbitrary network variation setup.
Remark 2 (Characteristic Variation and Exogenous Effects).
In our setup, we do not assume observation of any characteristics other than agent identity. Part of the reflection problem of Manski (1993) is disentangling exogenous and endogenous group effects. If we add characteristic variation, we can distinguish between exogenous and endogenous effects in a two step procedure. In the first step, we proceed with the identification arguments from Section 3.1.1. Upon doing so, we have recovered $v_{i}$ for each agent. In the second step, we treat $v_{i}$ as the outcome variable and proceed with standard analysis of the relationship between $v_{i}$ and observed characteristics. Notably, our first stage captures the endogenous effects of the reflection problem and the second stage captures the exogenous effects.
Remark 3 (Product-Attribute Variation).
Another variation that could be introduced in our framework is product attribute variations. Suppose that there is a set of $K$ observable attributes, with $a_{x}$ denoting the vector of attributes for product $x$. $a_{x}$ includes not only things that affect product quality, but also things like price, advertising, etc. We assume that the ideal point of each agent is a linear function of observable product attributes. This is captured by an agent-specific real vector $\beta_{i}$ such that $u_{i}(x)=\beta_{i}a_{x}$ for each $i$ and $x$. This means that we assume that attributes affect utility in the same way for all products. Then the ideal point is calculated as $v_{i}(x)=\frac{u_{i}(x)}{\sum_{y\in X}u_{i}(y)}$. Identification of $\beta_{i}$ then proceeds with a two-step procedure. As above, we first recover $v_{i}$ for each agent. In the second step, we treat $v_{i}$ as the outcome variable and proceed with standard analysis of the relationship between $v_{i}$ and observed attributes.
5 Related Literature
In this paper we study the testing and identification properties of the linear-in-means model when an analyst observes group variation. As part of our analysis, we show that the reflection problem of Manski (1993) is a generic problem when each agent’s decision space is one dimension but stops being generic when we move to higher dimensions. This highlights the identifying power of more granular data in the linear-in-means model and we emphasize this as a key takeaway of our analysis. We now conclude with a discussion of this paper’s place in the related literature.
Our paper is related to several strands of literature. We begin by discussing the strand which studies the linear-in-means model of social interactions. Predating the modern literature on the linear-in-means model, Keynes (1937) considers a model of financial markets via a story of beauty contests. In this setting, an agent wishes to take the action that coincides with the average action of the rest of the population. In our setup, this corresponds to $\pi_{i}^{N}(i)=0$ and $\pi_{i}^{N}(j)=\frac{1}{|N|-1}$. More recently, Ushchev and Zenou (2020) studies the microfoundations and comparative statics of the linear-in-means model allowing for arbitrary network structure. As mentioned earlier, Ushchev and Zenou (2020) along with Blume et al. (2015), Boucher and Fortin (2016), and Kline and Tamer (2020) show that the linear-in-means choice rule can be achieved as the best response to a quadratic loss utility function in a complete information game where each agent knows each $v_{i}$ and $\pi_{i}^{N}$. Golub and Morris (2020) consider an extension of this setup where agents have incomplete information and relate the linear-in-means model to higher-order expectations as well as conventions in networks. We build on this literature by studying the empirical content of the linear-in-means model.
More closely related to our paper is the strand of literature which focuses studying the identification properties of the linear-in-means model as well as other related models of peer effects. A seminal contribution in this literature is Manski (1993) and his discussion of the reflection problem. Important to our analysis is the following takeaway from the reflection problem of Manski (1993). It is in general difficult to identify the social influence of a group on an agent due to the endogenous nature of outcomes. In our setting, this corresponds to identifying both the underlying network structure and the corresponding weights on directed edges in this network. Much of the literature following Manski (1993) aims to identify social interaction parameters when the underlying network structure is (partially) known. This literature is extensive, so we list the following, all of which provide various conditions in order to recover identification of social interaction parameters in linear-in-means style models; Lee (2007), Graham (2008), Bramoullé et al. (2009), De Giorgi et al. (2010), Blume et al. (2011), Boucher et al. (2014), Blume et al. (2015), De Paula (2017). More recently, Boucher et al. (2024) extends the linear-in-means to a CES in means model of social influence and provides identifications in their setting.
Perhaps most closely related to our analysis within this literature is the work of Lewbel et al. (2023) and De Paula et al. (2024). Both of these papers study the joint identification of social interaction parameters as well as the identification of the underlying network structure. [^19] Breza et al. (2020) and Griffith and Peng (2023) also considers joint identification of peer effects and network structure. However, Breza et al. (2020) asks for a stronger form of data where the analyst knows the distribution over how many connection each agent has to other agents with specific characteristics. Similarly, Griffith and Peng (2023) makes assumptions on the underlying network density in order to recover the unobserved network structure. Battaglini et al. (2022) also estimates network connections in a related model of endogenous network formation. While Lewbel et al. (2023) utilizes group variation and De Paula et al. (2024) utilizes long run panel data, a common tool for identification in each of these studies is the use of characteristic variation in order to recover identification. The models of these two paper also differs in the fact that the characteristics of each agent show up in the choice procedure of other agents. Adapting to our context, the core components of these models can be summarized by
$p_{i}^{N}=\pi_{i}^{N}(i)v_{i}+\alpha z_{i}+\beta\sum_{j\in N\setminus i}\pi_{i}^{N}(j)p_{j}^{N}+\gamma\sum_{j\in N\setminus i}\pi_{i}^{N}(j)z_{j}$ (11)
where $z_{i}$ is the characteristic vector of agent $i$, $\alpha$ is the average impact of characteristics on each agent’s bliss point, $\beta$ corresponds to the average impact of other agents’ choice on each agent’s choice, and $\gamma$ corresponds to the average impact of each other agents’ characteristics on each agent’s choice. De Paula et al. (2024) interprets $\beta$ as the endogenous peer effects and $\gamma$ as the exogenous peer effects from the reflection problem of Manski (1993). Our analysis differs from this literature in that we require no knowledge of characteristics (other than identity) to jointly recover each agent’s bliss point as well as the underlying network structure. The key to our identification result is that our outcome vector $p_{i}^{N}$ is multidimensional. In fact, generically, our results show that identification fails in our setup when the outcome variable is one dimension.
Our paper also contributes to the choice theoretic literature studying social interactions. To our knowledge, this literature begins with Cuhadaroglu (2017) who studies a two period model with social influence taking effect in the second period. Borah et al. (2018) also considers a two period model of social influence. In the first period, social influence is used to form an agent’s consideration set and the second period is used for choice. Kashaev et al. (2023) consider a model of social influence where the choices of an agent’s peers directly form their consideration set. One of the main goals of this work is the actual identification of underlying social parameters, which they are able to achieve through dynamic choice variation. Bhushan et al. (2023) considers a model of choice where agents influence each other through their beliefs. An agent’s beliefs are determined through a process similar to ULM. Choices then correspond to subjective expected utility given these beliefs. Most closely related to our work in this literature is the work of Chambers et al. (2023) who consider a version of the linear-in-means model. They focus on a setting with menu variation (i.e. variation of $X$) with a fixed group and network structure. This differs from our analysis in that our focus is on group variation and we can accommodate network variation.
More generally, our paper is related to the literature studying the empirical content of strategic settings. Part of this literature focuses on testing the empirical content of specific solution concepts. Sprumont (2000) studies the testable content of Nash equilibrium. The work of Haile et al. (2008) finds that quantal response equilibrium has no empirical content. Similarly, Bossert and Sprumont (2013) and Rehbeck (2014) find that backwards induction has no empirical content when we only observe the induced choice function. Another portion of this literature focuses on characterizing the empirical content of specific (types of) games. To begin, Lee (2012) characterizes the testable content of zero-sum games. Carvajal et al. (2013) provides a revealed preference style characterization of the Cournot model of competition. Finally, Lazzati et al. (2025) characterizes the empirical content of Nash equilibrium in monotone games. As mentioned earlier, the linear-in-means model can be thought of as arising from Nash equilibrium play where each agent’s utility is given by Equation 2. As a result of this equivalence, we characterize Nash equilibrium play in the corresponding game when we observe player variation.
Finally, our paper is also related to the literature in stochastic choice studying agents who have preferences for non-deterministic bundles. The idea of deliberately stochastic preferences goes back to at least Machina (1985). Recently, Cerreia-Vioglio et al. (2019) axiomatizes data that arises from agents choosing with deterministic preferences over lotteries. Similarly, Fudenberg et al. (2015) characterizes stochastic choice data that arises from agents who have cardinal preferences over each alternative but face a perturbation to their utility function within the simplex. Allen and Rehbeck (2019) studies the identification properties of a similar class of perturbed utility functions. There has been little work on these types of perturbed utility functions with social influence components. Hashidate and Yoshihara (2023) is an exception and considers a perturbed utility function where the perturbation corresponds to a norm. Recall that the choices in the linear-in-means model can be induced by a perturbed utility model where an agent’s base utility is given by a quadratic loss function with reference to their ideal point. The perturbation then corresponds to a sum of quadratic loss functions with each one referencing another agent’s choice.
Appendix
Appendix A Preliminary Results
We begin with a preliminary lemma that formalizes our discussion at the start of Section 3.1.
Lemma A.1.
$p_{i}^{N}=\gamma_{i}v_{i}+\sum_{j\in N\setminus i}\gamma_{j}p_{J}^{N}$ for $v_{i}\in\Delta(X)$ with $\gamma_{j}\geq 0$, $\gamma_{i}>0$, and $\sum_{j\in N}\gamma_{j}=1$ if and only if $v_{i}\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Proof.
Suppose that $p_{i}=\gamma_{i}v_{i}+\sum_{j\in N\setminus i}\gamma_{j}p_{J}^{N}$ for $v_{i}\in\Delta(X)$ with $\gamma_{j}\geq 0$ and $\sum_{j\in N}\gamma_{j}=1$. Since $0<\gamma_{i}\leq 1$, we can recover that $v_{i}=\frac{1}{\gamma_{i}}p_{i}^{N}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}p_{j}^{N}$. Since $\gamma_{j}\geq 0$ and $\sum_{j\in N}\gamma_{j}=1$, $\frac{-\gamma_{j}}{\gamma_{i}}\leq 0$ and $\frac{1}{\gamma_{i}}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}=1$ and so $v_{i}\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Now suppose that $v_{i}\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$. This tells us that $v_{i}=\sum_{j\in N}\gamma_{j}p_{j}^{N}$ with $\gamma_{j}\leq 0$ $\forall j\in N\setminus i$ and $\sum_{j\in N}\gamma_{j}=1$. It follows that $\gamma_{i}\geq 1$. By basic algebra, we can solve for $p_{i}$ and get $p_{i}=\frac{1}{\gamma_{i}}v_{i}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}p_{j}^{N}$. Since $\gamma_{j}\geq 0$ for $j\in N\setminus i$ and $\sum_{j\in N}\gamma_{j}=1$, $\frac{-\gamma_{j}}{\gamma_{i}}\geq 0$ and $\frac{1}{\gamma_{i}}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}=1$ and so $p_{i}=\pi_{i}v_{i}+\sum_{j\in N\setminus i}\pi_{j}p_{J}^{N}$ for $v_{i}\in\Delta(X)$ with $\pi_{j}\geq 0$, $\pi_{i}>0$, and $\sum_{j\in N}\pi_{j}=1$. ∎
Define $X_{1}=\{x\in\mathbb{R}^{n}:\sum_{i}x_{i}=1\}$. So $X_{1}$ is not the simplex but rather the hyperplane containing the simplex. The following very simple result is closely related to classical results, for example Girsanov (1972), Lemma 3.11, or characterizations of Pareto optimality, relating to Negishi weights e.g. Smale (1974); Wan (1975) but is written to be close in form to Billot et al. (2000); Samet (1998). The main difference from the previous pair of papers is that no compactness assumptions are claimed, but rather polyhedrality of the sets is required.
Theorem 4.
Let $A_{1},\ldots,A_{k}$ each be nonempty subsets of $X_{1}$, and suppose additionally that each $A_{i}$ is a polyhedron. Suppose that $\bigcap_{i=1}^{k}A_{i}=\varnothing$. Then the following are equivalent:
  1. 1.
    $\bigcap_{i=1}^{k}A_{i}=\varnothing$.
  2. 2.
    For each $i=1,\ldots,k$, there are $p_{i}\in\mathbb{R}^{n}$ for which $\sum_{i=1}^{k}p_{i}=0$ and such that $p_{i}\cdot x_{i}>0$ for all $x_{i}\in A_{i}$.
Proof.
We show that 1 implies 2. Observe that $X_{1}$ is itself defined by a finite set of linear inequalities. We now consider $Y=\{(x,\ldots,x):x\in X_{1}\}\subseteq\mathbb{R}^{nk}$ and $\prod_{i=1}^{k}A_{i}$. These sets are both clearly polyhedra. Further, by hypothesis, $Y\cap\prod_{i=1}^{k}A_{i}=\varnothing$. By Corollary 19.3.3 and Theorem 11.1 of Rockafellar (1970), there exists $(p_{1},\ldots,p_{k})\in\mathbb{R}^{nk}$ and $c\in\mathbb{R}$ for which for all $x\in X_{1}$, $\sum_{i}p_{i}\cdot x<c$ and for all $(x_{1},\ldots,x_{k})\in\prod_{i}A_{i}$, $\sum_{i}p_{i}\cdot x_{i}>c$. In particular we may choose $c$ so that $\inf_{(x_{1},\ldots,x_{k})\in\prod_{i}A_{i}}\sum_{i}p_{i}\cdot x_{i}>c$.
Define $p_{i}^{*}=p_{i}-(c/k)\mathbf{1}$ and observe that, because all $A_{i}$ are subsets of $X_{1}$, the $c$ in the inequalities gets replaced with 00 when replacing $p_{i}$ with $p_{i}^{*}$. Specifically, this also tells us that (by considering indicator functions of the form $\mathbf{1}_{j}$ for $j\in\{1,\ldots,n\}$), we have $\sum_{i}p_{i}^{*}\ll 0$. Further, it tells us that for each $x,y\in X$, $\sum_{i}p_{i}^{*}(x)(k+1)-\sum_{i}p_{i}^{*}(y)k<0$, as $(k+1)\mathbf{1}_{x}-k\mathbf{1}_{y}\in X_{1}$. Thus $\sum_{i}\frac{(k+1)p_{i}^{*}(x)-kp_{i}^{*}(y)}{2k+1}<0$, and by taking limits, $\sum_{i}p_{i}^{*}(x)\leq\sum_{i}p_{i}^{*}(y)$. Since $x$ and $y$ are arbitrary, this implies that $\sum_{i}p_{i}^{*}(x)=\sum_{i}p_{i}^{*}(y)$ for all $x,y\in X$.
Now, define for each $i=1,\ldots k$, $c_{i}=\inf_{x_{i}\in A_{i}}p_{i}^{*}\cdot x_{i}$. Then we have $\sum_{i}c_{i}>0$. Fix $\epsilon>0$ so that $\sum_{i}(c_{i}-\epsilon)>0$. For each $i$ then let $q_{i}^{*}=p_{i}^{*}-(c_{i}-\epsilon)\mathbf{1}$ and observe that for all $x_{i}\in A_{i}$, $q_{i}^{*}\cdot x_{i}=p_{i}^{*}\cdot x_{i}-(c_{i}-\epsilon)\cdot x_{i}\geq\epsilon>0$ as $x_{i}\in X_{1}$. Further, $\sum_{i}q_{i}^{*}\ll 0$ remains valid as $\sum_{i}(c_{i}-\epsilon)>0$. Further, it remains true that $\sum_{i}q_{i}^{*}(x)=\sum_{i}q_{i}^{*}(y)$ for all $x,y\in X$ as we have simply subtracted a constant from each $q_{i}^{*}$.
We show that 2 implies 1. Suppose by means of contradiction that there is $x^{*}\in\bigcap_{i=1}^{k}A_{i}$. Then $p_{i}\cdot x^{*}>0$ for all $i$ and in particular then $(\sum_{i=1}^{k}p_{i})\cdot x^{*}>0$. But this is a contradiction as $(\sum_{i=1}^{k}p_{i})\cdot x^{*}=0\cdot x^{*}=0$. ∎
Appendix B Omitted Proofs
B.1 Proof of Theorem 1
Our proof of the equivalence between conditions (1) and (2) utilizes Lemma A.1 and follows quickly from it. Our proof of the equivalence between (1) and (3) utilizes Theorem 4. For the separation, let us consider the sets $\Delta(X)$ and, for each $N$, $\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})$ where
$\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})=\{v|v=\sum_{j\in N}\gamma_{j}p_{j}^{N},\gamma_{j}\leq 0\text{ }\forall j\in N\setminus i,\sum_{j\in N}\gamma_{j}=1\}.$ (12)
Note that this differs from Equation 4 as $v$ is not required to be in the simplex. Owing to Rockafellar (1970) Theorem 19.3, each $\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})$ is polyhedral.
Proof.
We begin with the equivalence between (1) and (2). Suppose that $\{p_{i}\}$ is data consistent with GLM. Then, for each $i\in\mathcal{A}$, there exist $v_{i}$ and $\pi_{i}^{N}$ for each $N\in\mathcal{N}_{i}$ such that Equation 1 holds for each $(i,N)$, Since $v_{i}$ is fixed across each $N\in\mathcal{N}$, by Lemma A.1, the collection of sets $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ must have a point of common intersection.
Now suppose that, for each $i\in\mathcal{A}$, the collection of sets $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ has a point of common intersection. By Lemma A.1, the $v$ in the intersection of $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ works as a feasible $v_{i}$ for each $N\in\mathcal{N}_{i}$. Specifically, for each $N\in\mathcal{N}_{i}$, there are $\gamma_{j}\geq 0$, with $\gamma_{i}>0$ and $\sum_{j\in N}\gamma_{j}=1$, such that $p_{i}^{N}=\gamma_{i}v+\sum_{j\in N\setminus i}\gamma_{j}p_{j}^{N}$. Take these $\gamma_{j}$ as $\pi_{i}^{N}(j)$. These $v$ and $\pi_{i}$ correspond to a GLM representation of the dataset.
Now we prove the equivalence between (1) and (3). By definition of no incentive compatible money pump is equivalent to the lack of existence of a bet satisfying strict feasibility, individual rationality, and incentive compatibility. This is then equivalent to the following. For every $i\in\mathcal{A}$, there does not exist a collection of vectors $\{b^{N}\}_{N\in\mathcal{N}_{i}}$ with $b^{N}\in\mathbb{R}^{X}$ and $\sum_{N\in\mathcal{N}_{i}}b^{N}\ll 0$ such that for all $N\in N_{i}$, $b^{N}\cdot p_{i}^{N}>0$ and for each $j\neq i$ and $N$ for which $j\in N$, $b^{N}\cdot(p_{i}^{N}-p_{j}^{N})\geq 0$. If there are $b^{N}$ as in the statement of the theorem, then the model is not valid. Suppose, by means of contradiction that there are such $b^{N}$ but the model is valid. Because $\pi_{i}^{N}(i)>0$, we know that $v_{i}=\frac{p_{i}^{N}}{\pi^{N}_{i}(i)}-\sum_{j\neq i}\frac{\pi_{i}^{N}(j)}{\pi_{i}^{N}(i)}p_{j}^{N}$. In particular, since $\sum_{x}v_{i}(x)=1$, we must have $\frac{1}{\pi^{N}_{i}(i)}-\sum_{j\neq i}\frac{\pi_{i}^{N}(j)}{\pi_{i}^{N}(i)}=1$. Let $\lambda_{i}=\frac{1}{\pi^{N}_{i}(i)}$ and $\lambda_{j}=\frac{\pi_{i}^{N}(j)}{\pi_{i}^{N}(i)}$, then we have $\lambda_{i}-\sum_{j\neq i}\lambda_{j}=1$, where each $\lambda_{k}\geq 0$. Then $\lambda_{i}p_{i}^{N}-\sum_{j\neq i}\lambda_{j}p_{j}^{N}=(\lambda_{i}-\sum_{j\neq i}\lambda_{j})p_{i}^{N}+\sum_{j\neq i}\lambda_{j}(p_{i}^{N}-p_{j}^{N})$. Consequently $b^{N}\cdot v_{i}=(\lambda_{i}-\sum_{j\neq i}\lambda_{j})b^{N}\cdot p_{i}^{N}+\sum_{j\neq i}\lambda_{j}b^{N}\cdot(p_{i}^{N}-p_{j}^{N})>0$. So for each $N$, we have $b^{N}\cdot v_{i}>0$. Consequently $(\sum_{N}b^{N})\cdot v_{i}>0$, but $\sum_{N}b^{N}\ll 0$, a contradiction.
On the other hand, suppose the model is violated. Recall $\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})$. Observe that every $\gamma$ in this set satisfies $\gamma_{i}>0$. Further, observe that $\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})$ reflects the set of possible $v_{i}$ which are compatible with the observed choice $p_{i}^{N}$. That is, if there were some $v_{i}$ common to all $\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})$, we could define $\pi_{i}^{N}(i)=\frac{1}{\gamma_{i}}$ and $\pi_{i}^{N}(j)=-\frac{\gamma_{j}}{\gamma_{i}}$. Now, by Theorem 19.3 of Rockafellar (1970), each $\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})$ is polyhedral (owing to the fact that the set of $\gamma_{j}:j\in N$ satisfying the linear inequalities is a polyhedron). Since the model is violated, there is no $v_{i}$ common to all $\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})$. Consequently $\Delta(X)\cap\bigcap_{N:i\in N}\hat{co}^{-1}(\Delta(p^{N}),p_{i}^{N})=\varnothing$. By Theorem 4, there are weights $b^{N}$ and $q$ for which $q+\sum_{N}b^{N}=0$, where $q\cdot x>0$ for all $x\in\Delta(X)$, which means $q\gg 0$, and where $b^{N}\cdot p_{i}^{N}>0$ (by taking a weight of one on $p_{i}^{N}$) and $b^{N}\cdot((k+1)p_{i}^{N}-kp_{j}^{N})>0$ for all $j$, which implies by taking limits with respect to $k$ and normalizing, $b^{N}\cdot(p_{i}^{N}-p_{j}^{N})\geq 0$. Finally, $\sum_{N}b^{N}=-q\ll 0$.
B.2 Proof of Corollary 1
Proof.
By Theorem 1 condition (2), we know that a dataset is consistent with GLM if and only if, for each $i$, $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}}$ have a point of common intersection. As mentioned in Section 3.1, if $p_{i}^{N}$ lies on the interior of $\Delta(p_{-i}^{N})$, then $N$ offers no testable content for agent $i$. Specifically, if $p_{i}^{N}$ lies on the interior of $\Delta(p_{-i}^{N})$ then $p_{i}^{N}$ can be written as a convex combination of $p_{-i}^{N}$ and $v\in\Delta(X)$ for any choice of $v$ with a vanishingly small weight put on $v$. It follows that we can restrict our attention to $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ for groups $N$ where $p_{i}^{N}\not\in\Delta(p_{-i}^{N})$. These groups are exactly $\mathcal{N}_{i}^{+}\cup\mathcal{N}_{i}^{-}$. We have a series of observations for three cases.
  1. 1.
    Suppose $N\in\mathcal{N}_{i}^{+}\setminus\mathcal{N}_{i}^{-}$
    In this case $p_{i}^{N}\geq p_{j}^{N}$ for all $j\in N\setminus i$ with a strict inequality for at least one $j$. Fix this $j$. In this case, set $\pi_{i}^{N}(k)=0$ for each $k\in N\setminus\{i,j\}$. It then follows that every $v$ satisfying $v\geq p_{i}^{N}$ can be rationalized by $p_{i}^{N}=\pi_{i}^{N}(i)v+\pi_{i}^{N}(j)p_{j}^{N}$ for some choice of convex weights, but any $v<p_{i}^{N}$ cannot.
  2. 2.
    Suppose $N\in\mathcal{N}_{i}^{-}\setminus\mathcal{N}_{i}^{+}$
    In this case $p_{i}^{N}\leq p_{j}^{N}$ for all $j\in N\setminus i$ with a strict inequality for at least one $j$. Fix this $j$. In this case, set $\pi_{i}^{N}(k)=0$ for each $k\in N\setminus\{i,j\}$. It then follows that every $v$ satisfying $v\leq p_{i}^{N}$ can be rationalized by $p_{i}^{N}=\pi_{i}^{N}(i)v+\pi_{i}^{N}(j)p_{j}^{N}$ for some choice of convex weights but any $v>p_{i}^{N}$ cannot.
  3. 3.
    Suppose $N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}$
    In this case $p_{i}^{N}=p_{j}^{N}$ for each $j\in N$. In this case, any convex combination of $v$ with $p_{j}^{N}$ for $j\in N\setminus i$ which puts positive weight on $v$ would be different from $p_{j}^{N}$ if and only $p_{i}^{N}\neq p_{j}^{N}$. Thus, in this case we have that $v_{i}=p_{j}^{N}=p_{i}^{N}$.
We now consider two cases.
  1. 1.
    Suppose that $\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}=\emptyset$.
    In this case $[0,\min_{N\in\mathcal{N}_{i}^{-}}p_{i}^{N}]=\bigcap_{N\in\mathcal{N}_{i}^{-}}\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}^{-}}$ and $[\max_{N\in\mathcal{N}_{i}^{+},1}p_{i}^{N},1]=\bigcap_{N\in\mathcal{N}_{i}^{+}}\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}^{+}}$. These two sets have a point of intersection if and only if $\max_{N\in\mathcal{N}_{i}^{+}}p_{i}^{N}\leq\min_{N\in\mathcal{N}_{i}^{-}}p_{i}^{N}$. The other two conditions are vacuous in this case.
  2. 2.
    Suppose that $\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}\neq\emptyset$.
    By observation 3 above, for each $N^{*}\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}$, $co^{-1}(\Delta(p^{N^{*}}),p_{i}^{N^{*}})$ is exactly $p_{i}^{N^{*}}$ and so it must be the case that $v_{i}=p_{i}^{N^{*}}=p_{i}^{=}$ if the dataset is consistent. It then follows that this value $v_{i}$ must be unique among consistent data. By the arguments from case 1, we need it to be that $\max_{N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}}p_{i}^{N}\leq\min_{N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}}p_{i}^{N}$. Further, it must also be the case that $\max_{N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}}p_{i}^{N}\leq p_{i}^{=}\leq\min_{N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}}p_{i}^{N}$. However, $N^{*}\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}$, so $\max_{N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}}p_{i}^{N}\geq p_{i}^{=}\geq\min_{N\in\mathcal{N}_{i}^{+}\cap\mathcal{N}_{i}^{-}}p_{i}^{N}$. It then follows that the three conditions hold if and only if we the dataset is consistent with GLM.
B.3 Proof of Proposition 1
Proof.
This follows immediately from Lemma A.1 and Theorem 1. ∎
B.4 Proof of Corollary 2
Proof.
By consistency with the linear-in-means model, $co^{-1}(\Delta(p^{N_{j}}),p_{i}^{N_{j}})$ and $co^{-1}(\Delta(p^{N_{k}}),p_{i}^{N_{k}})$ intersect. By $N_{j}$ and $N_{k}$ being binary, $co^{-1}(\Delta(p^{N_{j}}),p_{i}^{N_{j}})$ and $co^{-1}(\Delta(p^{N_{k}}),p_{i}^{N_{k}})$ are one-dimensional rays, and so they intersect once or infinitely often. By linear independence of $(p_{i}^{N_{j}}-p_{j}^{N_{j}})$ and $(p_{i}^{N_{k}}-p_{k}^{N_{k}})$ and by the definition of $co^{-1}(\Delta(p^{N_{j}}),p_{i}^{N_{j}})$ and $co^{-1}(\Delta(p^{N_{k}}),p_{i}^{N_{k}})$ , these two sets intersect at a single point. By Proposition 1, this single point of intersection is exactly our identified set for $v_{i}$. ∎
B.5 Proof of Proposition 2
Proof.
As $v_{i}$ and $\{p_{k}^{N}\}_{k\in N\setminus i}$ are affinely independent, any point in their convex hull can be written as a convex combination of these points with unique weights on each point. In the linear-in-means model, we have $p_{i}^{N}=\pi^{N}_{i}(i)v_{i}+\sum_{j\neq i}\pi^{N}_{i}(j)p^{N}_{j}$, and so $p_{i}^{N}$ is in the convex hull of these points and thus each $\pi^{N}_{i}(j)$ is uniquely pinned down. ∎
B.6 Proof of Theorem 2
Proof.
To begin observe that no weakly incentive compatible money pump is satisfied if and only if there does not exist a collection of vectors $\{\alpha^{N}\}_{N\in\mathcal{N}_{i}}$ with $\alpha^{N}\in\mathbb{R}^{X}$ and $\sum_{N\in\mathcal{N}_{i}}\alpha^{N}\ll 0$ such that for all $N\in\mathcal{N}_{i}$, $\alpha^{N}\cdot p_{i}^{N}>0$ and for each $j\neq i$, $\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}\cdot(p_{i}^{N}-p_{j}^{N})\geq 0$. Now suppose that LLM is satisfied, and by means of contradiction, that there exist $\alpha^{N}$ as in the previous sentence. Drop dependence of $w_{i}$ on $i$ to simplify notation. We use the notation $w(N)=\sum_{j\in N}w(j)$.
Observe that
$\sum_{N\in\mathcal{N}_{i}}w(N)\alpha^{N}\cdot p_{i}^{N}=$ $\sum_{N\in\mathcal{N}_{i}}\left(w(i)\alpha^{N}\cdot v_{i}+\sum_{j\in N\setminus i}w(j)\alpha^{N}\cdot p_{j}^{N}\right)$
$=$ $w(i)\left(\sum_{N\in\mathcal{N}_{i}}\alpha^{N}\right)\cdot v_{i}+\sum_{j\in\mathcal{A}\setminus i}w(j)\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}\cdot p_{j}^{N}$
$<$ $\sum_{j\in\mathcal{A}\setminus i}w(j)\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}\cdot p_{j}^{N}$
$\leq$ $\sum_{j\in\mathcal{A}\setminus i}w(j)\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}\cdot p_{i}^{N}$
$=$ $\sum_{N\in\mathcal{N}_{i}}\sum_{j\in N\setminus i}w(j)\alpha^{N}\cdot p_{i}^{N}.$
Here, the strict inequality follows as $v_{i}\in\Delta(X)$, $w(i)>0$, and $\sum_{N\in\mathcal{N}_{i}}\alpha^{N}\ll 0$. The weak inequality follows as $\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}(p_{i}^{N}-p_{j}^{N})\geq 0$ and $w(j)\geq 0$. The equalities are by definition and algebraic manipulation. Now, subtracting the right hand side of the string of inequalities from the left hand side, we obtain:
$\sum_{N\in\mathcal{N}_{i}}w(i)\alpha^{N}\cdot p_{i}^{N}<0.$
This contradicts the facts that $w(i)>0$ and for each $N\in\mathcal{N}_{i}$, $\alpha^{N}\cdot p_{i}^{N}>0$.
For the other direction, we want to find, for each $i$, numbers $f_{i}(x)\in\mathbb{R}$, and for each $j\neq i$, $w_{i}(j)\in\mathbb{R}$ and finally for each $N$ for which $i\in N$, $\lambda_{i}(N)$ such that the following equations are satisfied:
  1. 1.
    $f_{i}(x)\geq 0$ for all $x$
  2. 2.
    $0<\sum_{x}f_{i}(x)$
  3. 3.
    $w_{i}(j)\geq 0$ for all $j\neq i$
  4. 4.
    $f_{i}(x)+\lambda_{i}(N)p_{i}^{N}(x)+\sum_{j\in N\setminus i}w_{i}(j)p_{j}^{N}(x)=0$ for all $x$ and all $N$ for which $i\in N$
If we find such numbers, define the Luce weights for agent $i$ as $w(i)=\sum_{x}f_{i}(x)>0$, $w(j)=w_{i}(j)\geq 0$ and define $v_{i}(x)=\frac{f_{i}(x)}{\sum_{x}f_{i}(x)}$. Observe then that by equation 4:
$w(i)v_{i}(x)+\lambda_{i}(N)p_{i}^{N}(x)+\sum_{j\in N\setminus i}w(j)p_{j}^{N}(x)=0$, so that
$-\lambda_{i}(N)p_{i}^{N}(x)=w(i)v_{i}(x)+\sum_{j\in N\setminus i}w(j)p_{j}^{N}(x).$
Since we know $\sum_{x}v_{i}(x)=1$ (by definition), $\sum_{x}p_{i}^{N}(x)=1$ and $\sum_{x}p_{j}^{N}(x)=1$ (by assumption), it follows automatically that $-\lambda_{i}(N)=w(N)$ and we have shown that this is a linear formulation of Luce.
The dual of this system, a Theorem of the Alternative, for example Motzkin’s Theorem, (Mangasarian, 1994) p. 28 implies the existence of $\beta^{N}\in\mathbb{R}^{X}$ for each $N\in\mathcal{N}_{i}$ for which
  1. 1.
    For all $N\in\mathcal{N}$ for which $i\in N$, $\beta^{N}\cdot p_{i}^{N}=0$.
  2. 2.
    For all $j\in\mathcal{A}\setminus i$, $\sum_{N\in\mathcal{N}_{i}:j\in N}\beta^{N}\cdot p_{j}^{N}\leq 0$.
  3. 3.
    $\sum_{N\in\mathcal{N}_{i}}\beta^{N}\ll 0$.
Now, owing to finiteness of $X$, we know that there exists some $b<0$ such that for each $x\in X$, $\sum_{N\in\mathcal{N}_{i}}\beta^{N}(x)<b$. Define $\alpha^{N}=\beta^{N}-\frac{b}{|\mathcal{N}_{i}|}\mathbf{1}$. Observe that $\alpha^{N}\cdot p_{i}^{N}=\beta^{N}\cdot p_{i}^{N}-\frac{b}{|\mathcal{N}_{i}|}=\frac{-b}{|\mathcal{N}_{i}|}>0$. Observe that $\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}\cdot p_{j}^{N}\leq\frac{-b|\{N\in\mathcal{N}_{i}:j\in N\}|}{|\mathcal{N}_{i}|}=\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}\cdot p_{i}^{N}$, so that $\sum_{N\in\mathcal{N}_{i}:j\in N}\alpha^{N}\cdot(p_{i}^{N}-p_{j}^{N})\geq 0$. Finally, $\sum_{N\in\mathcal{N}_{i}}\alpha^{N}=(\sum_{N\in\mathcal{N}_{i}}\beta^{N})-b\mathbf{1}\ll 0$.
Appendix C GLM Results when $\pi_{i}^{N}(i)\geq 0$
In this section we consider an extension of GLM where we now allow $\pi_{i}^{N}(i)\geq 0$ (allowing for equality). We call this version of the model GLM. Our goal is to discuss and prove how our main results on GLM extend to GLM. Recall that for an agent $i$, $\mathcal{N}_{i}^{ext}$ corresponds to the set of groups $N$ containing $i$ with $p_{i}^{N}\not\in\Delta(p_{-i}^{N})$. Observe that in GLM, whenever $p_{i}^{N}$ lies in $\Delta(p_{-i}^{N})$, $p_{i}^{N}$ is consistent with GLM as it can be written as a convex combination of $p_{-i}^{N}$. This motivates the following.
Definition 12.
A set of vectors $\{b^{N}\}_{N\in\mathcal{N}_{i}^{ext}}$ with $b^{N}\in\mathbb{R}^{X}$ for each $N\in\mathcal{N}_{i}$ is called a ext bet on agent i.
Definition 13.
A bet on agent $i$ is ext strictly feasible if $\sum_{N\in\mathcal{N}_{i}^{ext}}b^{N}\ll 0$.
Definition 14.
A bet on agent $i$ is ext individually rational if $b^{N}\cdot p_{i}^{N}>0$ for each $N\in\mathcal{N}_{i}^{ext}$.
Definition 15.
A bet on agent $i$ is ext incentive compatible if $b^{N}\cdot(p_{i}^{N}-p_{j}^{N})\geq 0$ for each $N\in\mathcal{N}_{i}^{ext}$ and each $j\in N\setminus i$.
Definition 16.
We say that a dataset $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies no ext incentive compatible money pump if for each $i\in\mathcal{A}$ there are no ext strictly feasible, ext individually rational, and ext incentive compatible ext bets on agent $i$.
Theorem 5.
For a dataset $\{p^{N}\}_{N\in\mathcal{N}}$, the following are equivalent.
  1. 1.
    $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with GLM.
  2. 2.
    For every $i\in\mathcal{A}$, the collection of sets $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}^{ext}}$ has a point of mutual intersection.
  3. 3.
    $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies no ext incentive compatible money pump.
Before proving Theorem 5, we first prove a preliminary lemma.
Lemma C.1.
Suppose $p_{i}^{N}\not\in\Delta(p_{-i}^{N})$. Then $p_{i}^{N}=\gamma_{i}v_{i}+\sum_{j\in N\setminus i}\gamma_{j}p_{J}^{N}$ for $v_{i}\in\Delta(X)$ with $\gamma_{j}\geq 0$ and $\sum_{j\in N}\gamma_{j}=1$ if and only if $v_{i}\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Proof.
Suppose that $p_{i}=\gamma_{i}v_{i}+\sum_{j\in N\setminus i}\gamma_{j}p_{J}^{N}$ for $v_{i}\in\Delta(X)$ with $\gamma_{j}\geq 0$ and $\sum_{j\in N}\gamma_{j}=1$. Since $p_{i}^{N}\not\in\Delta(p_{-i}^{N})$, it must be the case that $\gamma_{i}>0$ and so $0<\gamma_{i}\leq 1$. By doing basic algebra we can recover that $v_{i}=\frac{1}{\gamma_{i}}p_{i}^{N}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}p_{j}^{N}$. Since $\gamma_{j}\geq 0$ and $\sum_{j\in N}\gamma_{j}=1$, $\frac{-\gamma_{j}}{\gamma_{i}}\leq 0$ and $\frac{1}{\gamma_{i}}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}=1$ and so $v_{i}\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Now suppose that $v_{i}\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$. This tells us that $v_{i}=\sum_{j\in N}\gamma_{j}p_{j}^{N}$ with $\gamma_{j}\leq 0$ $\forall j\in N\setminus i$ and $\sum_{j\in N}\gamma_{j}=1$. It follows that $\gamma_{i}\geq 1$. By basic algebra, we can solve for $p_{i}$ and get $p_{i}=\frac{1}{\gamma_{i}}v_{i}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}p_{j}^{N}$. Since $\gamma_{j}\geq 0$ for $j\in N\setminus i$ and $\sum_{j\in N}\gamma_{j}=1$, $\frac{-\gamma_{j}}{\gamma_{i}}\geq 0$ and $\frac{1}{\gamma_{i}}+\sum_{j\in N\setminus i}\frac{-\gamma_{j}}{\gamma_{i}}=1$ and so $p_{i}=\pi_{i}v_{i}+\sum_{j\in N\setminus i}\pi_{j}p_{J}^{N}$ for $v_{i}\in\Delta(X)$ with $\pi_{j}\geq 0$ and $\sum_{j\in N}\pi_{j}=1$. ∎
We now proceed to prove Theorem 5.
Proof.
The equivalence between (1) and (2) follows from Lemma C.1 and our discussion at the start of this section. To see the equivalence between (1) and (3), observe that, in the proof of Theorem 1, the arguments from the proof of (1$\implies$(3) go through replacing $\mathcal{N}_{i}$ with $\mathcal{N}_{I}^{ext}$ and observing that once $\pi_{i}^{N}\not\in\Delta(p_{-i}^{N})$ any coefficient preceding $\pi_{i}^{N}$ will be non-zero (see the first two sentences of the proof of Lemma C.1). To show (3) implies (1), observe that the arguments from (3$\implies$(1) from the proof of Theorem 1 hold when we try and prove the separation of $\Delta(X)$ and $\{co^{-1}(\Delta(p^{N}),p_{i}^{N})\}_{N\in\mathcal{N}_{i}^{ext}}$. ∎
Now let $\mathcal{N}_{i}^{-*}\subseteq\mathcal{N}_{i}$ denote the set of groups $N$ satisfying $p_{i}^{N}<p_{j}^{N}$ for each $j\in N\setminus i$. Similarly, let $\mathcal{N}_{i}^{+*}\subseteq\mathcal{N}_{i}$ denote the set of groups $N$ satisfying $p_{i}^{N}>p_{j}^{N}$ for each $j\in N\setminus i$.
Corollary 3.
In the one dimension case, a dataset $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with GLM if and only if, for all $i\in\mathcal{A}$, $\min_{N\in\mathcal{N}_{i}^{-*}}p_{i}^{N}\geq\max_{\mathcal{N}_{i}^{+*}}p_{i}^{N}$ when both $\mathcal{N}_{i}^{-*}$ and $\mathcal{N}_{i}^{+*}$ are non-empty.
Proof.
Observe that $\mathcal{N}_{i}^{-*}\cup\mathcal{N}_{i}^{+*}=\mathcal{N}_{i}^{ext}$ in the one dimension case. Further, by the arguments in the proof of Corollary 1, $co^{-1}(\Delta(p^{N}),p_{i}^{N})=[0,p_{i}^{N}]$ when $N\in\mathcal{N}_{i}^{-*}$ and $co^{-1}(\Delta(p^{N}),p_{i}^{N})=[p_{i}^{N},1]$ when $N\in\mathcal{N}_{i}^{+*}$. It then follows from Theorem 5 that our corollary holds. ∎
Proposition 4.
In GLM, the sharp identified set for $v_{i}$ is given by $\bigcap_{N\in\mathcal{N}_{i}^{ext}}co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Proof.
This is an immediate consequence of Lemma C.1 and Theorem 5. ∎
All of the other identification results for GLM carry through to GLM as written.
Appendix D Relation Between Theorem 1 and Samet (1998)
In this appendix, we discuss the relation between Samet (1998) and our characterizations of GLM and GLM. The main theorem of Samet (1998) provides a linear program that characterizes when a collection of compact convex sets fail to have a point of mutual intersection. In light of Theorems 1 and 5, the result of Samet (1998) can be used as an alternative characterization for GLM and GLM.
Theorem 6.
The following are equivalent.
  1. 1.
    $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with the general linear-in-means model.
  2. 2.
    For every $i\in\mathcal{A}$ there does not exist a collection of vectors $\{\alpha^{N}\}_{N\in\mathcal{N}_{i}}$ with $\alpha^{N}\in\mathbb{R}^{X}$ and $\sum_{N\in\mathcal{N}_{i}}\alpha^{N}=0$ such that, for all $N\in\mathcal{N}_{i}$, $\alpha^{N}\cdot v>0$ for each $v\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Theorem 6 immediately follows from condition (2) of Theorem 1, the main result of Samet (1998), and the observation that each $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ is compact and convex. Part of the original motivation for Samet (1998) was the no interim trade result of Morris (1994). In the setting of Samet (1998) and Morris (1994), $\{\alpha^{N}\}_{N\in\mathcal{N}_{i}}$ corresponds to a trading scheme, the equality condition corresponds to a feasibility condition, and the inequality condition corresponds to an individual rationality condition.
We now interpret Theorem 6 in the context of a no trade story. Suppose we have a collection of outside observers. We have one observer for each group $N$ in $\mathcal{N}_{i}$ and we index each observer by their corresponding group $N$. These observers are risk neutral but ambiguity averse (in the max-min sense of Gilboa and Schmeidler (1989)) and consider state contingent trades. In this setting, a state corresponds to an alternative $x$ and the frequency of state $x$ corresponds to the frequency with which agent $i$ would choose $x$ in isolation, $v_{i}(x)$. In this sense, $\{\alpha^{N}\}_{N\in\mathcal{N}_{i}}$ corresponds to a trade scheme in the language of Morris (1994) where $\alpha^{N}(x)$ corresponds to the payout to observer $N$ when agent $i$ chooses $x$ in isolation. Each observer knows that agent $i$’s choices follow from Equation 1. The information available to observer $N$ is $p^{N}$, the set of choices made in group $N$. From this, observer $N$ is able to form a set of beliefs about the choices of agent $i$. This set of beliefs corresponds to $co^{-1}(\Delta(p^{N}),p_{i}^{N})$. The condition $\sum_{N\in\mathcal{N}_{i}}\alpha^{N}=0$ corresponds to a trading scheme being balanced and $\alpha^{N}\cdot v>0$ for each $v\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$ corresponds to the trading scheme being profitable in expectation for every belief in observer $N$’s set of beliefs. With this interpretation, Theorem 6 says that there is no trade between these outside observers if and only if the dataset is consistent with GLM. In other words, there is no trade if and only if each agent’s ideal point $v_{i}$ is common across their groups.
Before moving on, we also note that condition (2) from Theorem 6 can be simplified somewhat. Since $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ is a convex set and $\alpha^{N}\cdot v>0$ is a linear constraint, instead of checking $\alpha^{N}\cdot v>0$ for each $v\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$, we can simply check $\alpha^{N}\cdot v>0$ for the extreme points of $co^{-1}(\Delta(p^{N}),p_{i}^{N})$. Since $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ is polyhedral, this is a finite process. Finally, we state the analogue of Theorem 6 for GLM.
Theorem 7.
The following are equivalent.
  1. 1.
    $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with GLM.
  2. 2.
    For every $i\in\mathcal{A}$ there does not exist a collection of vectors $\{\alpha^{N}\}_{N\in\mathcal{N}_{i}^{ext}}$ with $\alpha^{N}\in\mathbb{R}^{X}$ and $\sum_{N\in\mathcal{N}_{i}^{ext}}\alpha^{N}=0$ such that, for all $N\in\mathcal{N}_{i}^{ext}$, $\alpha^{N}\cdot v>0$ for each $v\in co^{-1}(\Delta(p^{N}),p_{i}^{N})$.
Theorem 7 immediately follows from condition (2) of Theorem 5, the main result of Samet (1998), and the observation that each $co^{-1}(\Delta(p^{N}),p_{i}^{N})$ is compact and convex.
Appendix E Proofs and Extension of Results from Section 3.3
In this section, we consider the ULM model of Section 3.3 and allow for variation in group size. With this in mind, we now introduce the extensions of the axioms from 3.3 allowing for group size variation.
Axiom 5 (Cyclically constant).
A dataset $\{p_{N}\}_{N\in\mathcal{N}}$ is cyclically constant if for every cycle $\{(i_{k},j_{k},N_{k})\}_{k=1}^{m}$ we have that $\sum_{k=1}^{m}(1+|N_{k}|)(p_{i_{k}}^{N_{k}}-p_{j_{k}}^{N_{k}})=0$.
Axiom 6 (Constant Peer Effects).
A dataset $\{p_{N}\}_{N\in\mathcal{N}}$ satisfies constant peer effects if for each $N$ and $M$ with $i,j\in N\cap M$ we have that $(1+|N|)[p_{i}^{N}-p_{j}^{N}]=(1+|M|)[p_{i}^{M}-p_{j}^{M}]$.
Axiom 7 (Symmetric Peer Effects).
For $i\in N\cap M$,
$p_{i}^{N}-\sum_{j\in N\setminus M}[p_{j}^{N}-p_{i}^{N}]=p_{i}^{M}-\sum_{k\in M\setminus N}[p_{j}^{M}-p_{i}^{M}]-\frac{|N|-|M|}{1+|N|}\sum_{l\in N\cap M}[p_{l}^{M}-p_{i}^{M}].$ (13)
Observe that in the previous three axioms, each axiom reduces to the axioms from the main text when each group has the same size.
Theorem 8.
For a dataset $\{p^{N}\}_{N\in\mathcal{N}}$, the following are equivalent.
  1. 1.
    $\{p^{N}\}_{N\in\mathcal{N}}$ is consistent with the uniform linear-in-means model.
  2. 2.
    $\{p^{N}\}_{N\in\mathcal{N}}$ satisfies cyclic constancy, symmetric peer effects, and bounded total peer effects.
Proof.
We begin by proving necessity of the three axioms. We begin with cyclic constancy. By the definition of ULM we have the following for $i,j\in N$.
$\begin{split}p_{i}^{N}-p_{j}^{N}&=\frac{1}{|N|}v_{i}+\sum_{k\neq i}\frac{1}{|N|}p^{N}_{k}-\frac{1}{|N|}v_{j}-\sum_{k\neq j}\frac{1}{|N|}p^{N}_{k}\\ &=\frac{1}{|N|}v_{i}-\frac{1}{|N|}v_{j}+\frac{1}{|N|}(p_{j}^{N}-p_{i}^{N})\\ &=\frac{1}{1+|N|}(v_{i}-v_{j})\end{split}$ (14)
The equivalence between the second and third line follows from the properties of geometric series. Now consider any cycle $\{(i_{k},j_{k},N_{k})\}_{k=1}^{m}$.
$\begin{split}\sum_{k=1}^{m}(1+|N_{k}|)(p_{i_{k}}^{N_{k}}-p_{j_{k}}^{N_{k}})&=\sum_{k=1}^{m}(1+|N_{k}|)(\frac{1}{1+|N_{k}|}(v_{i_{k}}-v_{j_{k}}))\\ &=\sum_{k=1}^{m}(v_{i_{k}}-v_{j_{k}})\\ &=0\end{split}$ (15)
The first line follows from substitution. The second line follows from canceling like terms. The third line follows from the fact that for cycles $j_{k}=i_{k+1}$, and $j_{m}=i_{1}$. To see that bounded total peer effects is necessary, observe the following.
$\begin{split}p_{i}^{N}&=\frac{1}{|N|}v_{i}+\frac{1}{|N|}\sum_{j\in N\setminus i}p_{j}^{N}\\ v_{i}&=|N|p_{i}^{N}-\sum_{j\in N\setminus i}p_{j}^{N}\\ &=p_{i}^{N}-\sum_{j\in N\setminus i}[p_{j}^{N}-p_{i}^{N}]\end{split}$ (16)
The first equality holds by definition of our model. The second and third equalities hold by rearrangement. Since we require our bliss point $v_{i}$ to be in the simplex, bounded total peer effect then holds by the last line. We now prove necessity of symmetric peer effects. By Equation 16 we have the following.
$\begin{split}p_{i}^{N}-\sum_{j\in N\setminus i}[p_{j}^{N}-p_{i}^{N}]=p_{i}^{M}-\sum_{k\in M\setminus i}[p_{k}^{M}-p_{i}^{M}]\\ p_{i}^{N}-\sum_{j\in N\setminus M}[p_{j}^{N}-p_{i}^{N}]+\sum_{l\in(N\cap M)\setminus i}\frac{1+|M|}{1+|N|}[p_{l}^{M}-p_{i}^{M}]\\ =p_{i}^{M}-\sum_{k\in M\setminus N}[p_{k}^{M}-p_{i}^{M}]-\sum_{l\in(N\cap M)\setminus i}[p_{l}^{M}-p_{i}^{M}]\\ p_{i}^{N}-\sum_{j\in N\setminus M}[p_{j}^{N}-p_{i}^{N}]=p_{i}^{M}-\sum_{k\in M\setminus N}[p_{j}^{M}-p_{i}^{M}]-\frac{|N|-|M|}{1+|N|}\sum_{l\in N\cap M}[p_{l}^{M}-p_{i}^{M}]\end{split}$ (17)
The first equality comes from Equation 16. The second equality comes from cyclic constancy (which we have proved holds). [^20] Note that here we only need to use binary cycles (i.e. $[i,j]->[j,i]$). The third equality follows from gathering like terms. Thus all three of our axioms are necessary.
We now proceed to show sufficiency of our three axioms. To do so we need to construct a $v_{i}\in\Delta(X)$ for each $i$ such that these $v_{i}$ are consistent with the data we observe. To begin, for each $i$, choose some $N$ such that $i\in N$. Define $\hat{v}_{i}$ as follows.
$\hat{v}_{i}=p_{i}^{N}-\sum_{j\in N\setminus i}(p_{j}^{N}-p_{i}^{N})$
By bounded total peer effect, $\hat{v}_{i}$ is non-negative. Since $p_{i}^{N}$ and $p_{j}^{N}$ are probabilities, $\sum_{x\in X}\hat{v}_{i}(x)=1$ and thus $v_{i}$ lies in the simplex. We now verify if our construction induces our observed data. Consider an arbitrary $M$ such that $i\in M$. We start with $\hat{p}_{i}^{M}$ as the choice probabilities implied by our construction.
$\begin{split}\hat{p}_{i}^{M}&=\frac{1}{|M|}\hat{v}_{i}+\frac{1}{|M|}\sum_{j\in M\setminus i}p_{j}^{M}\\ &=\frac{1}{|M|}(p_{i}^{N}-\sum_{j\in N\setminus i}[p_{j}^{N}-p_{i}^{N}]+\sum_{j\in M\setminus i}p_{j}^{M})\\ \hat{p}_{i}^{M}-p_{i}^{M}&=\frac{1}{|M|}(p_{i}^{N}-\sum_{j\in N\setminus i}[p_{j}^{N}-p_{i}^{N}]-(p_{i}^{M}-\sum_{j\in M\setminus i}[p_{j}^{M}-p_{i}^{M}]))\\ &=\frac{1}{|M|}(p_{i}^{N}-\sum_{j\in N\setminus M}[p_{j}^{N}-p_{i}^{N}]+\\ &\sum_{l\in(N\cap M)\setminus i}\frac{1+|M|}{1+|N|}[p_{l}^{M}-p_{i}^{M}]-(p_{i}^{M}-\sum_{j\in M\setminus i}[p_{j}^{M}-p_{i}^{M}]))\\ &=\frac{1}{|M|}(p_{i}^{N}-\sum_{j\in N\setminus M}[p_{j}^{N}-p_{i}^{N}]-(p_{i}^{M}-\sum_{j\in M\setminus N}[p_{j}^{M}-p_{i}^{M}])\\ &-\frac{|N|-|M|}{1+|N|}\sum_{l\in N\cap M}[p_{l}^{M}-p_{i}^{M}])\\ &=0\end{split}$ (18)
The first equality follows from the definition of our model. The second equality follows from our definition of $\hat{v}_{i}$. The third equality follows by subtracting $p_{i}^{M}$ from both sides. The fourth equality follows from cyclic constancy. [^21] Again, here we only use binary cycles. The fifth equality follows from collecting like terms. The sixth equality follows from symmetric peer effects. We have constructed $\hat{v}_{i}$ which is in the simplex and induces our observed data. Thus cyclic constancy, symmetric peer effects, and bounded total peer effects are sufficient. ∎
Proposition 5.
Cyclic constancy holds iff for each $i\in\mathcal{A}$, there is $v_{i}$ for which $\sum_{x}v_{i}(x)=1$ and for each $N\in\mathcal{N}$, $O^{N}\in\mathbb{R}^{X}$ such that $\sum_{x}O^{N}(x)=0$ such that for all $N$ and $i\in N$, $p_{i}^{N}(x)=\frac{1}{|N|}v_{i}(x)+\frac{1}{|N|}\sum_{j\in N\setminus i}p_{j}^{N}(x)+O^{N}(x)$.
Remark 4.
In Proposition 5, we have not specified conditions on $v_{i}$ and $O^{N}$ ensuring that $p_{i}^{N}$ is actually a probability measure.
Proof.
Fix any $i^{*}\in\mathcal{A}$ and define $v_{i^{*}}$ arbitrarily so that $\sum_{x}v_{i^{*}}=1$; we do not require that $v_{i^{*}}\geq 0$. Define $v_{j}=3(p_{j}^{\{i^{*},j\}}-p_{i^{*}}^{\{i^{*},j\}})$. Cyclic constancy implies that for any $N\in\mathcal{N}$ and $\{i,j\}\subseteq N$, $(1+|N|)(p_{i}^{N}-p_{j}^{N})=v_{i}-v_{j}$. That is, $(1+|N|)(p_{i}^{N}-p_{j}^{N})=3(p_{i}^{\{i,i^{*}\}}-p_{i^{*}}^{\{i,i^{*}\}})-3(p_{j}^{\{j,i^{*}\}}-p_{i^{*}}^{\{j,i^{*}\}})=v_{i}-v_{j}$. Define the uniform Luce rule according to Luce weights $v_{i}$ (this may not necessarily have probabilities as frequencies here). Call this rule $U_{i}^{N}$, so that for each $N\in\mathcal{N}$ and each $i\in N$, $U_{i}^{N}=\frac{1}{|N|}v_{i}+\sum_{j\in N\setminus i}U_{j}^{N}$. By definition, for all $N$ and all $i,j\in N$ with $i\neq j$, we have: $(p_{i}^{N}-p_{j}^{N})=(U_{i}^{N}-U_{j}^{N})$ (this equality obviously holds when $i=j$). Consequently for all such $i,N$:, $p_{i}^{N}=U_{i}^{N}+(p_{j}^{N}-U_{j}^{N})$.
By taking averages $p_{i}^{N}=U_{i}^{N}+\frac{1}{|N|}\sum_{j\in N}(p_{j}^{N}-U_{j}^{N})$. Now let $\overline{O}^{N}=\frac{1}{|N|}\sum_{j\in N}(p_{j}^{N}-U_{j}^{N})$. Observe that $\sum_{x}\overline{O}^{N}(x)=0$. Therefore, for all $N\in\mathcal{N}$ and all $i\in N$, $p_{i}^{N}=U_{i}^{N}+\overline{O}^{N}$. Finally, using the representation of $U_{i}^{N}$, we obtain $p_{i}^{N}=\frac{1}{|N|}v_{i}+\frac{1}{|N|}\sum_{j\in N\setminus i}U_{j}^{N}+\overline{O}^{N}$, and using the fact that $U_{j}^{N}=p_{j}^{N}-\overline{O}^{N}$, we get
$p_{i}^{N}=\frac{1}{|N|}v_{i}+\frac{1}{|N|}\sum_{j\in N\setminus i}(p_{j}^{N}-\overline{O}^{N})+\overline{O}^{N}$
or $p_{i}^{N}=\frac{1}{|N|}v_{i}+\frac{1}{|N|}\sum_{j\in N\setminus i}p_{j}^{N}+\frac{1}{|N|}\overline{O}^{N}$. Setting $O^{N}=\frac{1}{|N|}\overline{O}^{N}$ establishes the result.
Conversely if we start with $v_{i}$ for each $i\in\mathcal{A}$ for which $\sum_{x}v_{i}(x)=1$ (not necessarily a probability), and for each $N\in\mathcal{N}$, a vector $O^{N}$ for which $\sum_{x}O^{N}(x)=0$, then if $p_{i}^{N}=U_{i}^{N}+O^{N}$ is a probability measure for all $N\in\mathcal{N}$ and $i\in N$, then simple algebra establishes that for any $N\in\mathcal{N}$ and $i,j\in N$, $(1+|N|)(p_{i}^{N}-p_{j}^{N})=v_{i}-v_{j}$. As a telescoping series, a sum across a cycle results in 00. ∎